### v1

# Trajectory Volatility for Out-of-Distribution Detection in Mathematical Reasoning

Abstract
Real-world data deviating from the independent and identically distributed (i.i.d.) assumption of in-distribution training data poses security threats to deep networks, thus advancing out-of-distribution (OOD) detection algorithms. Detection methods in generative language models (GLMs) mainly focus on uncertainty estimation and embedding distance measurement, with the latter proven to be most effective in traditional linguistic tasks like summarization and translation. However, another complex generative scenario mathematical reasoning poses significant challenges to embedding-based methods due to its high-density feature of output spaces, but this feature causes larger discrepancies in the embedding shift trajectory between different samples in latent spaces. Hence, we propose a trajectory-based method TV score, which uses trajectory volatility for OOD detection in mathematical reasoning. Experiments show that our method outperforms all traditional algorithms on GLMs under mathematical reasoning scenarios and can be extended to more applications with high-density features in output spaces, such as multiple-choice questions.
Figure 1: Illustration of the comparisons of input/output space between mathematical reasoning and text generation scenarios. We select the MATH [7] dataset for mathematical reasoning and the OPUS [52] for text generation (translation). We use SimCSE [11] to obtain sentence embeddings and project them via UMAP [38]. Different colors represent different domains, and lighter and darker shades of the same color represent the input and output spaces, respectively. We also present input/output examples of two domains in each scenario. Appendix B.1 and B.2 show detailed settings and full examples. Projections show that input/output space overlapping of different domains under the reasoning scenario is more obvious, especially in output spaces where high-density collapse happens.
1 Introduction
The rapid development of generative language models (GLMs) [43, 44, 6, 2, 53] has empowered them to fit diverse and challenging datasets, showing strong generalization over in-distribution (ID) test data satisfying the independent and identically distributed (i.i.d.) assumption. However, unconstrained inputs in real-world settings frequently trigger distributional drifts, called out-of-distribution (OOD) data. In such scenarios, model performance often deteriorates unexpectedly, yielding harmful outcomes. Thus, OOD detection [41, 4] is critical in safeguarding model security.
Based on the availability of OOD data, detection methods can be classified into two categories. When OOD data is available, constructing a binary classifier with ID and OOD classes [9, 24, 19] is a standard method. However, the OOD data variety in real-world settings makes it impractical for classifiers to learn all OOD data features, causing scalability challenges. Thus, we focus on more generalized scenarios where OOD data is unavailable, and methods can be categorized into three main types: (1) Output-based methods assess confidence by analyzing predicted probabilities [14, 5, 34, 55, 30]; (2) Ensemble-based methods assess the uncertainty through a collection of supporting models [10, 22]; (3) Feature-based methods assess the Mahalanobis Distance between OOD samples and ID data distribution in the feature space [26, 45, 51]. Besides, there are some fragmented methods, such as gradient methods [27, 17] and auto-encoder reconstruction methods [8, 18]. However, these fragmented methods encounter optimization and computational complexity issues, leading to significant performance bottlenecks [58], so they receive relatively less attention.
Most of the above methods were proposed for classification tasks, while the research on deploying OOD detection methods on GLMs is relatively niche. Existing methods on GLMs focus on uncertainty estimation [37, 35] and embedding distance measure [46], and [46] has demonstrated that embedding-based methods are currently the only optimal solution for text generation scenarios such as summarization and translation. However, embedding-based methods encounter challenges under mathematical reasoning scenarios, as shown in Figure 1: (1) Input Space: In the text generation scenarios, input questions from different domains show more obvious clustering features, yet that under mathematical reasoning are not well distinguished. We will delve deeper into this issue in Appendix G.2; (2) Output Space: The output space shows high density under mathematical reasoning, we call this phenomenon “pattern collapse”. This is because the output space of mathematical reasoning is scalar, resulting in a compressed search space with a higher probability of overlap in two widely divergent questions. For example, “$1+3=$” and “$\int_{1}^{3}x\mathrm{d}x=$” are both “$4$”. Therefore, current methods suffer serious performance bottlenecks in mathematical reasoning scenarios.
Therefore, we transform our perspective from the static embedding representation to the dynamic embedding shift trajectory in the latent space and make the following findings: (1) Due to the pattern collapse property in the output space, trajectories corresponding to different samples are more likely to differentiate due to the hard constraints on the endpoints of trajectories compared to text generation scenarios. This theoretical perspective will be presented in Sec.2.1; (2) There exists early stabilization of GLMs for mathematical reasoning. For ID samples, GLMs have completed the main reasoning at a later stage, but this is not the case for the OOD samples. This empirical perspective will be presented in Sec.2.2. These indicate that trajectory is a promising tool to distinguish ID and OOD data.
Based on these analyses, we propose the Trajectory Volatility detection algorithm (TV score) for mathematical reasoning. We conduct experiments on mainstream mathematical reasoning datasets and GLMs for the basic OOD detection, and also consider the OOD quality estimation scenario, which imposes more fine-grained requirements on the OOD scores. Results show that our method outperforms all traditional algorithms on GLMs under the mathematical reasoning scenario.
Overall, our contribution is progressively three-fold:
  • We find phenomena of GLMs under mathematical reasoning about embedding low discrepancy in the input space, pattern collapse in the output space, and early stabilization in the latent space.
  • Based on the excavated phenomena, we verify the rationality of using trajectory as OOD detection from empirical and theoretical perspectives and propose a novel detection algorithm TV Score, which distinguishes ID and OOD samples by the embedding shift trajectory volatility.
  • Our method significantly outperforms traditional algorithms on GLMs under mathematical reasoning in diverse evaluation scenarios, including offline detection, online discrimination, and quality estimation, and it can be extended to more applications like multiple-choice questions.
1.1 Problem Statement
Mathematical reasoning relies on generative language models (GLMs). Let $\mathcal{X}$ be the input space and $\mathcal{Y}$ be the output space. $P_{\mathcal{X},\mathcal{Y}}$ is the joint data distribution defined over $\mathcal{X}\times\mathcal{Y}$, $P_{\mathcal{X}}$ and $P_{\mathcal{Y}}$ are the marginal distributions for inputs and outputs, respectively. GLMs are trained given input sequence ${\boldsymbol{x}}=x_{1}x_{2}...x_{t}\sim P_{\mathcal{X}}$ of length $t$ to autoregressively generate the next token in the corresponding output sequence ${\boldsymbol{y}}=y_{1}y_{2}...y_{n}\sim P_{\mathcal{Y}}$ of length $n$ over the likelihood model $p_{\boldsymbol{\theta}}({\boldsymbol{y}}|{\boldsymbol{x}})=\prod_{i=1}^{n}p_{\boldsymbol{\theta}}(y_{i}|{\boldsymbol{y}_{\prec i}},{\boldsymbol{x}})$, where each $x_{i}$ and $y_{i}$ are $d$-dimensional vectors and $\boldsymbol{\theta}$ is sampled from the parameter space $\Theta$.
Assume that $\widetilde{P}_{\mathcal{X},\mathcal{Y}}$ denote a distribution sufficiently different from $P_{\mathcal{X},\mathcal{Y}}$, the goal of OOD detection in GLMs is to find a score function $f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})$ for each sample and a threshold $\epsilon$, which may rely on the features of $\mathcal{X}$, $\mathcal{Y}$, and $\Theta$, to achieve a high discrimination accuracy goal:
$\max_{f}~{}\mathbb{P}_{(\boldsymbol{x},\boldsymbol{y})\sim P_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})<\epsilon\right]+\mathbb{P}_{(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}})\sim\widetilde{P}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}},\boldsymbol{\theta})>\epsilon\right],$ (1)
and a minimum expected risk where for any two different abnormal distributions $\widetilde{P}^{\prime}_{\mathcal{X},\mathcal{Y}}$ and $\widetilde{P}^{\prime\prime}_{\mathcal{X},\mathcal{Y}}$:
$\min_{f}\left|\mathbb{P}_{(\boldsymbol{x},\boldsymbol{y})\sim\widetilde{P}^{\prime}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\theta})>\epsilon\right]-\mathbb{P}_{(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}})\sim\widetilde{P}^{\prime\prime}_{\mathcal{X},\mathcal{Y}}}\left[f(\boldsymbol{\widetilde{x}},\boldsymbol{\widetilde{y}},\boldsymbol{\theta})>\epsilon\right]\right|.$ (2)
This is a robust goal but tough to achieve in practice, so the first goal is all current work’s main goal.
1.2 Notation and Definition
For a given sample $s_{i}$, $[{\boldsymbol{y}_{i}}]_{l}\in\mathbb{R}^{d}(0\leq l\leq L)$ denote the hidden embedding of the $l$-th layer, where $d$ is the embedding dimension and ${\boldsymbol{y}_{i}}]_{l}=\left([y_{i1}]_{l},[y_{i2}]_{l},...,[y_{id}]_{l}\right)^{\top}$. We set the embedding coordinate of the input space to $(0,[{\boldsymbol{y}_{i}}]_{0})$, the output space to (i.e., the last hidden layer) $(L,[{\boldsymbol{y}_{i}}]_{L})$, and the $L-1$ hidden layers to $\{(l,[{\boldsymbol{y}_{i}}]_{l})\}_{l=1}^{L-1}$.
Definition 1 (Component-independent Trajectory Volatility)
For sample $s_{i}$, its Component-independent Trajectory Volatility during all $L$ hidden layers is defined as the embedding difference between all neighboring layers:
${\boldsymbol{V}}(s_{i})=\frac{1}{L-1}\sum_{l=1}^{L-1}\left(|[y_{i1}]_{l+1}-[y_{i1}]_{l}|,|[y_{i2}]_{l+1}-[y_{i2}]_{l}|,...,|[y_{id}]_{l+1}-[y_{id}]_{l}|\right)^{\top}.$ (3)
Higher values of the $d$-th component of ${\boldsymbol{V}}(s_{i})$ indicate increased instability of the $d$-th dimension in the inference process for the sample $s_{i}$.
2 Motivation: Why Trajectory As The Measure?
2.1 Likelihood-High Trajectory Variation: A Macro Theoretical Perspective
Figure 2: The “pattern collapse” phenomenon only exists in mathematical reasoning scenarios, where two samples initially distant in distance will converge approximately at the endpoint after undergoing embedding shifts, and does not occur in text generation scenarios. This produces a greater likelihood of trajectory variation under different samples in mathematical reasoning.
Figure 1 has illustrated the phenomenon of pattern collapse within the output space in mathematical reasoning scenarios, and the scalar nature of the output space makes the phenomenon intuitive. However, this special phenomenon makes detecting OOD with trajectories in mathematical reasoning scenarios more promising. We abstract this phenomenon in Figure 2, which compares the trajectory trend between different samples in the mathematical reasoning and text generation scenarios.
In mathematical reasoning, when two initial points are separated by any distance in the input space, they typically converge to a significantly closer distance in the output space after undergoing an embedding shift. However, in text generation, outputs from different samples may not exhibit this same convergence. Constraints on trajectory endpoints in mathematical reasoning allow for a greater likelihood of variation in trajectory volatility under different samples. Based on this intuition, we propose a key hypothesis:
Hypothesis 1
In scenarios characterized by pattern collapse in the output space, the probability of variations in the volatility of embedding shift trajectories across samples increases.
This hypothesis reflects that we are more likely to distinguish between ID and OOD samples by the trajectory volatility in mathematical reasoning than in text generation. We specify that $[{\boldsymbol{y}}_{i}]_{L}\sim\mathcal{N}(\boldsymbol{c},\Sigma^{2})$. According to the pattern collapse properties under different tasks, we can constraint the endpoint embedding $[{\boldsymbol{y}}_{i}]_{L}$ in the output space:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\boldsymbol{O}$, so we approximate that $[{\boldsymbol{y}}_{i}]_{L}\equiv\boldsymbol{c}$;
[…]
With such constraints, we model our key hypothesis as a main theorem:
Theorem 2.1 (Main Theorem)
We assume that $\{[{\boldsymbol{y}}_{i}]_{l}\}_{l=1}^{L}$ are all independent variables sampling from $d-$dimemsional vector space $\mathbb{R}^{d}$. For different samples $s_{i}$ and $s_{j}$, the likelihood of variations in trajectory volatility under mathematical reasoning scenarios is higher than that under text generation scenarios, which means:
$\mathbb{E}_{\{[{\boldsymbol{y}}_{i}]_{l}\}_{l=1}^{L},~{}\{[{\boldsymbol{y}}_{j}]_{l}\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})\neq{\boldsymbol{0}}|[{\boldsymbol{y}}_{i}]_{L},[{\boldsymbol{y}}_{j}]_{L}\equiv\boldsymbol{c}\right\}$
[…]
Due to space limits, We present the complete Theoretical Proof in Appendix A. This proof can demonstrate that learning trajectories are prone to diverge between distinct samples when the output space experiences pattern collapse. This divergence increases the likelihood of discerning samples with varying features during reasoning tasks, as observed from the perspective of learning trajectories.
2.2 Early Stabilization: A Micro Empirical Perspective
Figure 3: Change curves of embedding differences between neighboring hidden layers after ID and OOD data are passed into GLMs (Llama-7B in this figure).
Figure 1 shows the minimal discrepancy in mathematical reasoning data’s input and output representations across various domains. This poses a significant risk to the efficacy of the current optimal embedding-based detection methods. However, these methods solely focus on the static nature of the embedding, neglecting its dynamic evolution within the latent space.
From this point, we first visualize the L1-norm of difference $\|\boldsymbol{y}_{l}-\boldsymbol{y}_{l-1}\|_{1}$ between the output $\boldsymbol{y}_{l}$ of each hidden layer and the output $\boldsymbol{y}_{l-1}$ of the previous layer under Llama-7B (32 layers), after ID and OOD data are input into the model, separately. Figure 3 shows the change curves within all hidden layers. We find that the embedding changes slightly until the 20th layer, and then enters the critical inference process after 20 layers. However, for ID data, the magnitude of embedding change is again suppressed after a few layers of inference, while for OOD data, the magnitude is maintained at a high level.
We refer to this phenomenon as Early Stabilization under mathematical reasoning: For ID data, GLMs largely complete their reasoning in the mid-to-late stages, and simple adjustments are sufficient after that. However, for OOD data, GLMs can still not complete accurate reasoning at a later stage. They thus can only randomly switch to a specific output pattern, i.e., the scalar immediate number pattern. This enables an intuition that leverages the trajectory for OOD detection may be effective.
3 TV Score: Trajectory Volatility for OOD Detection
In this section, we propose our trajectory-based algorithm TV Score to detect OOD samples in mathematical reasoning scenarios. Formally, for each input sample sequence ${\boldsymbol{x}}=x_{1}x_{2}...x_{n}$, the hidden embedding of $i$-th layer corresponding to the $i$-th token $x_{i}$ is denoted as ${\boldsymbol{h}_{i}^{l}}\in\mathbb{R}^{d}$. We set $n_{l}$ as the sequence length in $l$-th layer, and ${\boldsymbol{y}_{l}}$ to be the average embedding in the $l$-th layer, namely
${\boldsymbol{y}_{l}}=\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}{\boldsymbol{h}_{i}^{l}}~{}~{}~{}~{}(1\leq i\leq n_{l},1\leq l\leq L).$ (4)
Next, we need to obtain the trajectory volatility $V_{l}$ in each layer, then computer the final TV Score as:
$\textrm{TV}:=\frac{1}{L-1}\sum_{l=1}^{L-1}V_{l}.$ (5)
Therefore, defining the trajectory volatility $V_{l}$ stands as a key concern. The volatility between each neighboring layer-pair can be intuitively defined by the difference between the two embedding vectors, so the L1-norm $\|\boldsymbol{y}_{l}-\boldsymbol{y}_{l-1}\|_{1}$ is the simplest way. However, anomalous changes in each dimension under vector volatility will greatly affect the measure of volatility, introducing potentially uncontrollable factors. To address the challenge, we want to find the most critical principal component that maps the vector in low dimensions.
We consider using Mahalanobis Distance (MaDis) [26] to complete this mapping. We follow the assumptions of previous work [3, 46] and fit all average embedding ${\boldsymbol{y}_{l}}$ of ID samples in each layer to a Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l},{\boldsymbol{\Sigma}}_{l})$. For a new sample, We measure the MaDis between its hidden embedding ${\boldsymbol{y}}_{l}$ of $i$-th layer and $\mathcal{G}_{l}$, the conversion function $f({\boldsymbol{y}_{l}}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ is
$f(\boldsymbol{y}_{l})=\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)~{}~{}~{}~{}(1\leq l\leq L)$ (6)
The underlying assumption is that when the MaDis between embeddings and corresponding Gaussian distribution of ID data in one layer and that of the neighboring layer has a larger difference, the embedding shift between two layers is more volatile. In this case, $V_{l}$ is defined as:
$V_{l}=\left|f(\boldsymbol{y}_{l})-f(\boldsymbol{y}_{l-1})\right|.$ (7)
Differential Smoothing (DiSmo).
Outliers in the trajectory may significantly impact feature extraction [25, 32] then lead to errors. To mitigate this, we explore higher-order differential smoothing techniques to enhance trajectory smoothness. We first define the Gaussian distribution difference and embedding difference based on the additivity of the Gaussian distribution as
$\mathcal{N}({\boldsymbol{\mu}}_{l}^{(k)},{\boldsymbol{\Sigma}}_{l}^{(k)})=\Delta^{(k)}\mathcal{G}_{l}$ $=\Delta^{(k-1)}\mathcal{G}_{l+1}-\Delta^{(k-1)}\mathcal{G}_{l}=\Delta^{(k-2)}\mathcal{G}_{l+2}-2\cdot\Delta^{(k-2)}\mathcal{G}_{l+1}+\Delta^{(k-2)}\mathcal{G}_{l}$ (8)
$=\cdots=\mathcal{N}(\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\boldsymbol{\mu}}_{l+k},~{}\sum_{i=0}^{k}\mathrm{C}_{k}^{i}{\boldsymbol{\Sigma}}_{l+k}),$
${\boldsymbol{y}_{l}^{(k)}}={\boldsymbol{y}_{l+1}^{(k-1)}}-{\boldsymbol{y}_{l}^{(k-1)}}=\cdots=\sum_{i=0}^{k}(-1)^{k+i}\mathrm{C}_{k}^{i}{\boldsymbol{y}}_{l+k},$ (9)
where $\mathrm{C}_{k}^{i}=\frac{k!}{i!(k-i)!}$. Similarly, we computer $k$-th order embedding difference MaDis:
$f^{(k)}(\boldsymbol{y}_{l})=\left[{\boldsymbol{y}_{l}^{(k)}}-{\boldsymbol{\mu}_{l}^{(k)}}\right]^{\top}\left[\Sigma_{l}^{(k)}\right]^{-1}\left[{\boldsymbol{y}_{l}^{(k)}}-{\boldsymbol{\mu}_{l}^{(k)}}\right],$ (10)
and in this case, $V_{l}$ is defined as:
$V_{l}=\left|f^{(k)}(\boldsymbol{y}_{l})-f^{(k)}(\boldsymbol{y}_{l-1})\right|.$ (11)
The complete pseudo-code is presented in Appendix C. In experiments, we choose $k\leq 5$ because if $k$ is too large, the over-smoothing phenomenon may happen and effective features will be eliminated.
4 Experiments
4.1 Setup
Dataset Selection.
For the ID dataset, we use the MultiArith [47], which consists of Math Word Problems on arithmetic reasoning. For the OOD datasets, we intuitively introduce two types of detection scenarios following [46]: (i) Far-shift OOD setting, we select the MATH [13] dataset with five domains of algebra, geometry, counting and probability, number theory, and precalculus; (ii) Near-shift OOD setting, we select five independent datasets: GSM8K [7], SVAMP [42], AddSub [15], SingleEq [21], and SingleOp [21]. We consider the ID data negative(-) and the OOD data positive(+). Refer to Appendix D.1 for basic information and OOD features of these datasets.
Data Split and Sampling.
Given the limited data size of MultiArith, totaling only 600 samples and lacking a standard division, we allocate 360 samples for training and 240 for testing. However, with such a small test set, randomness in evaluation becomes a concern. To mitigate this, we conduct test sampling and set the sampling size as 1000. Specifically, we denote ID dataset as $\mathcal{D}_{\text{in}}$ and OOD dataset as $\mathcal{D}_{\text{out}}$. For each sampling, the collection is $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$ where $\widetilde{\mathcal{D}}_{\text{out}}\subset\mathcal{D}_{\text{out}}$ and $|\mathcal{D}_{\text{in}}|=|\widetilde{\mathcal{D}}_{\text{out}}|$, this guarantees positive and negative sample balance. We report both the mean and standard variance of the results to enhance the reliability of evaluations. Refer to Appendix D.2 for the ID dataset split.
Implementation.
To measure the application value of our method used in cutting-edge GLMs, we use Llama2-7B [53] and GPT-2-XL (1.5B) [6] as our backbones for ID dataset training. Refer to Appendix D.3 for training details. However, there exists uncertainty about the data used in the pre-training phase, especially for Llama2 because its data is closed-source. Some research [57, 59] have confirmed the absence of data leakage in Llama2 for the MATH and GSM8K datasets, we still conduct pre-experiments to examine the rationality of the OOD data selection rigorously. Our criterion is the claim that a dataset can be categorized as OOD if it exceeds the capabilities of the base model, as proposed in prior studies [50, 31]. Results of the pre-experiments are shown in Appendix D.4, and they can confirm that these datasets can be considered as OOD data for the two GLMs.
Baseline.
We compare our method with some training-free baselines [^1] We refer to the latest survey [23] to select baselines for the scarcity of OOD detection methods on GLMs. where OOD training data are unavailable: (1) Maximum Softmax Probability (MS-Prob) [14]; (2) Monte-Carlo Dropout (MC-Drop) [10]; (3) Sequence Perplexity (PPL) [3]; (4) Input Embedding (I-Emb) [46]; (5) Output Embedding (O-Emb) [46]. Additionally, we set the smoothing order $k$ ranges from 1 to 5 and report the highest among them when with smoothing. Refer to Appendix D.5 for baseline details.
Evaluation.
We establish two main evaluation scenarios: (i) The basic OOD detection scenario, we further divide it into offline and online discriminations; (ii) More challenging and fine-grained OOD quality estimation. Task-specific metrics are presented in the corresponding result sub-sections.
4.2 Scenario I: OOD Detection
Offline Discrimination.
We utilize TV scores for offline OOD discrimination. For each collection $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$, we report the AUROC [39] and FPR95 metrics. The former represents the area under the ROC curve and the latter represents the value of FPR at 95% TPR. Table 1 presents the results.
  • Performance Analysis: In the far-shift OOD setting, our average performance surpasses 98 (Llama2-7B) and 96 (GPT2-XL) under the AUROC metric, surpassing the optimal baseline by 10+ points. Moreover, our performance stands at an impressive 5.21 (Llama2-7B) and 9.89 (GPT2-XL) under the FPR95 metric, representing a remarkable 80%+ reduction compared to the optimal baseline, far surpassing all baseline methods. In the near-shift OOD setting, the robustness of our method is even more impressive. All of the baseline methods show significant performance degradation, especially in Llama2-7B, with the AUROC metric decreasing below 60 and the FPR95 elevating above 80. However, our method maintains excellent performances, with AUROC scores surpassing 90 and FPR95 below 30. This indicates that for more fine-grained OOD detection scenarios, our method demonstrates greater adaptability.
  • Model Analysis: Comparing performances of Llama2-7B and GPT2-XL, we find two phenomena: (i) Results on GPT2-XL are more stable, performance differences between GPT2-XL on far- and near-drift settings are not significant, while Llama2-7B shows a significant performance degradation (mainly for baselines) on near-shift setting; (ii) the DiSmo technique is more effective on GPT2-XL, which suggests that there are more anomalous learning tendencies in latent spaces of small models, and the smoothing helps to minimize these anomalies.
In addition, we conduct significant tests (Details are shown in Table 8 - 11). We find that our methods almost pass all significance tests, while the embedding-based methods have the lowest pass rate among baselines, suggesting that their results are more susceptible to sampling error.
Model Llama2-7B [53] GPT2-XL [6]
Method Metric Far-shift OOD Near-shift OOD Far-shift OOD Near-shift OOD
AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$
MS-Prob [14] 78.66$\pm$1.38 81.44$\pm$3.56 60.14$\pm$1.54 88.91$\pm$2.41 70.54$\pm$1.42 78.29$\pm$2.02 67.12$\pm$1.20 76.27$\pm$2.66
MC-Drop [10] 68.63$\pm$2.21 87.04$\pm$4.88 52.33$\pm$2.21 91.92$\pm$1.89 66.18$\pm$1.87 84.69$\pm$1.65 63.54$\pm$1.72 78.08$\pm$2.50
PPL [3] 85.64$\pm$1.46 53.06$\pm$4.36 59.35$\pm$1.89 86.09$\pm$1.89 80.82$\pm$1.04 64.53$\pm$2.10 73.74$\pm$1.12 72.39$\pm$1.27
I-Emb [46] 75.89$\pm$1.03 67.87$\pm$3.69 60.33$\pm$1.37 84.65$\pm$2.53 86.26$\pm$0.84 49.33$\pm$2.10 83.22$\pm$0.88 52.90$\pm$3.16
O-Emb [46] 74.86$\pm$1.39 75.21$\pm$2.16 44.50$\pm$1.06 86.46$\pm$1.59 77.95$\pm$1.16 65.64$\pm$3.42 79.28$\pm$1.24 64.70$\pm$2.72
TV score (Ours) 98.76$\pm$0.11 5.21$\pm$0.98 92.64$\pm$0.39 28.39$\pm$1.38 93.47$\pm$0.08 24.10$\pm$0.95 94.86$\pm$0.23 13.82$\pm$0.36
w/ DiSmo (Ours) 93.25$\pm$0.76 41.82$\pm$4.69 56.99$\pm$1.41 88.01$\pm$1.71 96.54$\pm$0.11 9.89$\pm$0.61 94.19$\pm$0.25 13.66$\pm$0.69
$\Delta$ (bold - underline) +13.12 -47.85 +32.31 -56.26 +10.28 -39.44 +11.64 -39.24
Table 1: OOD Detection — Offline Discrimination: AUROC and FPR95 results. Underline and bold denote SOTA among all baselines and all methods, respectively. We report the average results under each setting in the main text, results of each dataset are shown in Table 8 and 9 (Appendix E).
Online Discrimination.
In this part, we utilize the TV score for online OOD discrimination. For each collection $\{\mathcal{D}_{\text{in}},\widetilde{\mathcal{D}}_{\text{out}}\}$, we obtain a detector and computer the optimal cut-off $\tau_{i}$ of Youden Index, which is at the point in the AUROC curve where $\mathrm{TPR}-\mathrm{FPR}$ is maximum. Then for all OOD samples $s\in\mathcal{D}_{\text{out}}-\widetilde{\mathcal{D}}_{\text{out}}$, we donate $t$ as the sampling size and computer the discrimination accuracy:
$\text{Accuracy}=\frac{1}{t}\sum_{i=1}^{t}\frac{\sum_{s\in\mathcal{D}_{\text{out}}-\widetilde{\mathcal{D}}_{\text{out}}}\mathbb{I}\left[\text{TV-Score}(s)\geq\tau_{i}\right]}{\left|\mathcal{D}_{\text{out}}-\widetilde{\mathcal{D}}_{\text{out}}\right|}.$ (12)
In addition, the discrimination accuracy should vary less under different data collections, reflecting the discriminator’s robustness. Therefore, we denote the Robustness metric as sampling variance.
Far-shift OOD Setting Near-shift OOD Setting
[…]
I-Emb. / O-Emb. / TV (ours) I-Emb. / O-Emb. / TV (ours)
Algebra 76.43 / 45.42 / 93.88 5.27 / 6.94 / 0.97 GSM8K 81.49 / 75.32 / 93.39 10.08 / 3.36 / 2.05
Geometry 74.32 / 54.79 / 94.47 2.44 / 2.43 / 1.65 SVAMP 68.66 / 63.33 / 94.88 5.26 / 3.54 / 2.13
Cnt.&Prob 50.31 / 27.55 / 93.74 9.99 / 2.34 / 2.36 AddSub 79.16 / 78.09 / 74.11 3.21 / 6.98 / 2.77
Num.Theory 85.80 / 54.38 / 92.08 3.31 / 11.45 / 2.34 SingleEq 59.83 / 72.56 / 93.15 11.57 / 3.14 / 3.17
Precalculus 80.33 / 88.50 / 99.28 6.13 / 1.38 / 0.67 SingleOp 69.38 / 62.20 / 95.75 4.00 / 2.37 / 2.45
Average 73.44 / 54.13 / 94.69 5.43 / 4.91 / 1.60 Average 71.70 / 70.30 / 90.26 6.82 / 3.88 / 2.51
Table 2: OOD Detection — Online Discrimination: Accuracy and Robustness results. We mainly compare our method with embedding-based methods, and bold denotes the best among these methods.
Table 2 presents the results in Llama2-7B. Compared to the embedding-based methods, our TV score obtains about an average of 20-point accuracy improvement in both far-shift OOD and near-shift OOD settings, and on some datasets, such as Cnt.&Prob, our TV score achieves more than 40 points of improvement. These all imply that TV Score can perform online discrimination of OOD samples more accurately. In addition, our TV score also possesses stronger robustness, which means that in real scenarios, we can find the optimal threshold more consistently in the face of different accessible ID and OOD data, reducing the potential riskiness due to uncontrollable data acquisition.
4.3 Scenario II: OOD Quality Estimation
In this part, we utilize the TV score for generative quality estimation (QE). For text generation, the QE performance is usually measured by calculating the correlation coefficient between automatic scores and human ratings. However, QE in mathematical reasoning scenarios is not a well-defined problem. For one mathematical question, its answer is either right or wrong, the intermediate state does not exist. For example, when the correct answer is $12.5$, it is difficult to judge which is better between generated answers of $1.25$ and $13.6$. The human approach may be to judge by comparing the difference value or similarity like Rouge [28] and BertScore [60] between the generated answer and the correct answer, which is unfair to the machine because there is a lot of randomness in the intermediary process of computation, and the solution pattern of machines is case-based [16], so it is not suitable to judge the machine-generated results with customized mathematical rules.
Therefore, we use the binary direct matching [^2] Although direct matching is the most accurate solution, it suffers from two issues: (i) Generated answers may include much noisy content, increasing the matching difficulty; (ii) Performances on GLMs of mathematical reasoning is poor, which unbalances the positive and negative samples and increases the randomness. to compare the model-generated answers with the correct answers. Considering the open-ended output of the GLMs, we give a loose matching condition, i.e., as long as the correct answer is included in the generated answer by the model, the generated answer is recognized as correct and the matching score is 1, otherwise the matching score is 0. We compute the Kendall rank correlation coefficient $\tau$ [48] and Spearman rank correlation coefficient [40] between each OOD score and the matching score.
Model Llama2-7B [53] GPT2-XL [6]
Method Metric Far-shift OOD Near-shift OOD Far-shift OOD Near-shift OOD
Kendall $\uparrow$ Spearman $\uparrow$ Kendall $\uparrow$ Spearman $\uparrow$ Kendall $\uparrow$ Spearman $\uparrow$ Kendall $\uparrow$ Spearman $\uparrow$
MS-Prob 0.024$\pm$0.020 0.038$\pm$0.020 0.038$\pm$0.018 0.026$\pm$0.018 0.066$\pm$0.015 0.044$\pm$0.016 0.057$\pm$0.018 0.057$\pm$0.022
PPL 0.050$\pm$0.015 0.045$\pm$0.016 0.074$\pm$0.017 0.050$\pm$0.018 0.036$\pm$0.014 0.038$\pm$0.017 0.035$\pm$0.018 0.058$\pm$0.019
I-Emb 0.078$\pm$0.016 0.102$\pm$0.017 0.036$\pm$0.018 0.115$\pm$0.017 0.059$\pm$0.012 0.098$\pm$0.016 0.012$\pm$0.018 0.068$\pm$0.016
O-Emb 0.058$\pm$0.018 0.025$\pm$0.017 0.038$\pm$0.015 0.012$\pm$0.017 0.050$\pm$0.012 0.016$\pm$0.017 0.036$\pm$0.017 0.029$\pm$0.021
TV score (Ours) 0.161$\pm$0.012 0.147$\pm$0.015 0.159$\pm$0.017 0.158$\pm$0.017 0.138$\pm$0.010 0.123$\pm$0.013 0.131$\pm$0.015 0.146$\pm$0.015
w/ DiSmo (Ours) 0.111$\pm$0.016 0.152$\pm$0.015 0.113$\pm$0.018 0.134$\pm$0.017 0.139$\pm$0.009 0.141$\pm$0.014 0.123$\pm$0.014 0.154$\pm$0.016
$\Delta$ (bold - underline) +0.083 +0.050 +0.085 +0.043 +0.073 +0.043 +0.074 +0.086
Table 3: OOD Quality Estimation: Kendall’s $\tau$ and Spearman correlation between various OOD scores and benchmark quality metric binary matching. Each column shows the correlation when ID and OOD samples are merged. Underline denotes the SOTA among all baselines, and bold denotes the SOTA among our methods. We report the average results under each setting in the main text, results of each dataset are shown in Table 10 and 11 (Appendix E).
Table 3 presents the results. For Llama2-7B, when compared with Kendall correlation, the correlation improvement of TV scores over SOTA baselines reaches up to 100% under both far-shift and near-shift OOD settings. Compared with Spearman correlation, TV scores demonstrate a correlation enhancement over SOTA baselines by up to 100% under far-shift OOD setting and 30% under near-shift OOD setting. GPT2-XL also demonstrates excellent performance. These findings indicate that our TV scores not only facilitate the binary discrimination of ID and OOD samples but also substantially reflect the quality and precision of generated mathematical reasoning.
4.4 Beyond Mathematical Reasoning
Apart from mathematical reasoning, our method also has a wider range of potential applications that can be extended to any task where the output space exhibits the pattern collapse property. An example would be multiple-choice questions, which is a popular evaluation tool in the era of large language models and also display the pattern collapse property due to the limited output space being confined to the “ABCD” four options. To verify the generalizability of our method, we conduct experiments using the multiple-choice dataset MMLU [12], and our method also outperforms all traditional algorithms in this setting. Results and analyses are shown in Appendix F.
5 Discussion
Hyperparameter: Smoothing Order $k$.
In the main experiments, we find that the performance of differential smoothing fluctuates greatly in different settings. Therefore, we conduct the ablation of the smoothing order $k$. Results and analyses are shown in Appendix G.1.
Dilemmas of Embedding Representations in Reasoning.
Figure 4: A toy embedding projection. $\mathrm{A}$ is the benchmark, $\mathrm{X}$ is the ID sample, $\mathrm{Y}$ and $\mathrm{Z}$ are OOD samples.
We explore why the input-embedding-based method is not powerful under mathematical reasoning, as prior work [46] has shown it to be very effective in text generation scenarios.
[…]
We also use powerful sentence embedding techniques like SimCSE [11] to verify this conclusion. Results and analyses are shown in Appendix G.2.
Can Chain-of-Thought Address Pattern Collapse?
An intuitive approach to addressing pattern collapse in the output space is to use chain-of-thought (CoT) techniques [54, 20] to expand the output space size. Therefore, we include the solution steps in the output to re-evaluate the output embedding results and find that CoT cannot address this issue. Results and analyses are shown in Appendix G.3.
6 Related Work
Data-unavailable OOD Detection: Generic Methods.
In scenarios where OOD data is unavailable. Detection methods are categorized into three main types: (1) Output-based methods assess confidence using predicted probabilities. (2) Ensemble-based methods assess the uncertainty of a collection of supporting models, classical techniques are Monte-Carlo dropout [49, 10] based on Bayesian inference and deep ensemble [22, 35]; (3) Feature-based methods assess the Mahalanobis Distance between OOD samples and the distribution of the ID data feature space, usually considering the input and output spaces [26, 45, 51], occasionally extending to specific hidden layer spaces [1]. Besides, there are some fragmented methods, such as gradient methods [27, 17] and autoencoder reconstruction methods [8, 18]. Still, these methods suffer from optimization and computational complexity with serious performance bottlenecks [58] and thus are not mainstream detection methods.
OOD Detection in GLMs.
Relatively few studies have explored OOD detection in GLMs, mainly in semantic parsing [35, 29], speech recognition [37], machine translation [56, 36], summarization [46], and they do not jump out of the frameworks of uncertainty estimation, ensemble-based methods, and embedding-based methods. To our knowledge, we are the first to study OOD detection on mathematical reasoning, and we have found the failure of traditional algorithms in this scenario. Mathematical reasoning is an important and difficult research topic in the era of LLMs, and this research is valuable for the scenario expansion of OOD detection algorithms on language models.
7 Conclusion
We propose the TV score, a lightweight OOD detection method for mathematical reasoning, which distinguishes between ID and OOD samples by the embedding trajectory volatility in the latent space. We identify bottlenecks in OOD detection for mathematical reasoning and prove them empirically and theoretically. Experiments show that our method substantially outperforms all traditional algorithms, and can be extended to more application scenarios beyond mathematical reasoning.
Limitations
In this part, we discuss the limitations of this paper and outline our efforts to make up for them, along with potential avenues for future exploration.
  • Dataset Size: Due to the difficulty of collecting and labeling mathematical reasoning data, dataset sizes in this field are generally small, mostly in the hundreds or thousands, making it difficult to obtain millions of data for training and reasoning as in translation or summarization tasks. To address this, we adopt test sampling to reduce the randomness under small-scale testing and mitigate the data imbalance.
  • Generalizability: Our motivation and excavated phenomena are the mathematical reasoning scenario, and thus our main experiments are mainly centered on mathematical reasoning. However, the motivation for the method is primarily related to pattern collapse in the output space, and thus our method can be theoretically extended to any task where the output space has this property, such as multiple-choice questions. We have provided a preliminary demonstration of the extension feasibility in Appendix F, and more exploration can be left to future work.
Ethics Statement
The data and models used in this work are sourced from the official version of the original paper, and we strictly adhere to the provided usage protocol. Regarding the data, no modifications have been made to the original dataset. Regarding the models, supervised fine-tuning and OOD data inference are involved. To mitigate the risk of uncontrollable outputs, all generated outputs in the experiments have been reviewed to ensure their safety. Furthermore, as our focus is solely on mathematical reasoning and does not involve sensitive content, we would not cause any potential societal impact.
Appendix A Motivation I: Theoretical Analysis
In this section, we theoretically analyze Why pattern collapse in the output space leads to a greater likelihood of volatility differences in trajectories under different samples, which corresponds to Hypothesis 1 in the main paper.
A.1 Problem Setup
Latent Space Embedding.
For a given sample $s_{i}$, $[{\boldsymbol{y}_{i}}]_{l}\in\mathbb{R}^{d}(1\leq l\leq L)$ denote the hidden embedding of the $l$-th layer. We define the embedding coordinate in the latent space as $\{(l,[{\boldsymbol{y}_{i}}]_{l})\}_{l=1}^{L}$.
Embedding Interpolation.
We assume the ${\boldsymbol{F}}_{i}(x)=\left[F_{i1}(x),F_{i2}(x),...,F_{id}(x)\right]^{\top}:\mathbb{R}\rightarrow\mathbb{R}^{d}$ with $d$ independent components fits the $L$ coordinates, representing a continuous learning trajectory for $s_{i}$, so $[{\boldsymbol{y}_{i}}]_{l}={\boldsymbol{F}}_{i}(l)$. ${\boldsymbol{F}}_{i}(x)$ is taken from the functional space $\boldsymbol{X}$, so $\{{\boldsymbol{F}}_{i}(l)\}_{l=1}^{L}$ are all independent variables. We constrain each component function of this ${\boldsymbol{F}}_{i}(x)$ to satisfy the $m$-order $(m\geq 3)$ derivability property. Under this setting, the definition of component-independent trajectory volatility (Eq. 3) equates to
${\boldsymbol{V}}(s_{i})$ $=\left[V_{1}(s_{i}),V_{2}(s_{i}),...,V_{L}(s_{i})\right]^{\top}$ (13)
$=\frac{1}{L}\sum_{l=1}^{L-1}\left[|F_{i1}(l)-F_{i1}(l-1)|,|F_{i2}(l)-F_{i2}(l-1)|,...,|F_{id}(l)-F_{id}(l-1)|\right]^{\top}$
Modeling.
We observe the “pattern collapse” phenomenon in the output space in Figure 1. We abstract this phenomenon in Figure 2, which compares the trajectory trend between different samples in the mathematical reasoning and text generation scenarios.
We specify that $[{\boldsymbol{y}}_{i}]_{L}={\boldsymbol{F}}_{i}(L)\sim\mathcal{N}(\boldsymbol{c},\Sigma^{2})$, where
$\boldsymbol{c}=[c_{1},c_{2},...,c_{d}]^{\top},~{}~{}~{}\Sigma=\mathrm{diag}(\delta_{1},\delta_{2},\cdots,\delta_{d}).$ (14)
According to the pattern collapse property under different tasks, we can constraint the endpoint embedding ${\boldsymbol{F}}_{i}(L)$ in the output space:
  • For mathematical reasoning with pattern collapse, $\Sigma\rightarrow\boldsymbol{O}$, so we approximate that ${\boldsymbol{F}}_{i}(L)\equiv\boldsymbol{c}$;
[…]
With such constraints, we model the main theorem:
Theorem A.1 (Main Theorem)
For different samples $s_{i}$ and $s_{j}$, the likelihood of variations in trajectory volatility under mathematical reasoning scenarios is higher than that under text generation scenarios, which means:
$\mathbb{E}_{\{{\boldsymbol{F}}_{i}(l)\}_{l=1}^{L},~{}\{{\boldsymbol{F}}_{j}(l)\}_{l=1}^{L}~{}\sim~{}\mathcal{U}(\mathbb{R}^{d})}\left\{{\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})\neq{\boldsymbol{0}}|{\boldsymbol{F}}_{i}(L),{\boldsymbol{F}}_{j}(L)\equiv\boldsymbol{c}\right\}$
[…]
where $\mathcal{U}(\mathbb{R}^{d})$ denotes the uniform distribution defined in $d$-dimensional real number space.
A.2 Prelinimary
Next, we move on to the formal proofs. We begin with some propositions and lemmas that will be useful in the main theorem.
Proposition A.1 (Lagrange Remainder Term)
In the Taylor Expansion expression
$f(x)=f(a)+\frac{{\rm d}f}{{\rm d}x}(x-a)+\frac{1}{2!}\cdot\frac{{\rm d}^{2}f}{{\rm d}x^{2}}(x-a)^{2}+\cdots+\frac{1}{n!}\cdot\frac{{\rm d}^{n}f}{{\rm d}x^{n}}(x-a)^{n}+R_{n+1}(x),$
the remainder $R_{n+1}(x)$ has the following property:
$\left|R_{n+1}(x)\right|=\left|\frac{1}{(n+1)!}\cdot\frac{{\rm d}^{n+1}f}{{\rm d}c^{n+1}}\cdot(x-a)^{n+1}\right|\leq\frac{M}{(n+1)!}\cdot\left|(x-a)^{n+1}\right|,$
where $c\in[a,x]$ or $c\in[x,a]$, and $M=\sup\left\{\left|\frac{{\rm d}^{n+1}f}{{\rm d}\xi^{n+1}}\right|:\xi\in[a,x]~{}\textrm{or}~{}\xi\in[x,a]\right\}>0$.
Proof.  We consider the case of $a<x$ with $a>x$ identically. We use the Fundamental Theorem of Calculus (FTC) for the most basic expansion of $f(x)$:
$f(x)=f(a)+\int_{a}^{x}\frac{{\rm d}f}{{\rm d}x_{1}}~{}{\rm d}x_{1}.$ (15)
Continue to use the FTC to expand the derivatives in integrals:
$f(x)$ $=f(a)+\int_{a}^{x}\frac{{\rm d}f}{{\rm d}x_{1}}~{}{\rm d}x_{1}=f(a)+\int_{a}^{x}\left(\frac{{\rm d}f}{{\rm d}a}+\int_{a}^{x_{1}}\frac{{\rm d^{2}}f}{{\rm d}x_{2}^{2}}~{}{\rm d}x_{2}\right)~{}{\rm d}x_{1}$ (16)
$=f(a)+\frac{{\rm d}f}{{\rm d}x}(x-a)+\int_{a}^{x}\int_{a}^{x_{1}}\frac{{\rm d^{2}}f}{{\rm d}x_{2}^{2}}~{}{\rm d}x_{2}~{}{\rm d}x_{1}$
=$\cdots$= ⋯
$=\sum_{k=0}^{n}\frac{1}{k!}\cdot\frac{{\rm d^{k}}f}{{\rm d}a^{k}}\cdot(x-a)^{k}+\int_{a}^{x}\int_{a}^{x_{1}}\int_{a}^{x_{2}}\cdots\int_{a}^{x_{n}}\frac{{\rm d^{n+1}}f}{{\rm d}x_{n+1}^{n+1}}~{}{\rm d}x_{n+1}~{}{\rm d}x_{n}\cdots~{}{\rm d}x_{1}.$
Therefore, the generalized remainder is known as
$R_{n+1}(x)$ $=\int_{a}^{x}\int_{a}^{x_{1}}\int_{a}^{x_{2}}\cdots\int_{a}^{x_{n}}\frac{{\rm d^{n+1}}f}{{\rm d}x_{n+1}^{n+1}}~{}{\rm d}x_{n+1}~{}{\rm d}x_{n}\cdots~{}{\rm d}x_{1}$ (17)
$=\int_{a}^{x}\frac{1}{n!}\cdot\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}\cdot(x-t)^{n}~{}{\rm d}t.$
We let $m_{n+1}=\min_{t\in[a,x]}\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}$ and $M_{n+1}=\max_{t\in[a,x]}\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}$, so
$m_{n+1}\int_{a}^{x}(x-t)^{n}~{}{\rm d}t\leq\int_{a}^{x}\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}\cdot(x-t)^{n}~{}{\rm d}t\leq M_{n+1}\int_{a}^{x}(x-t)^{n}~{}{\rm d}t$ (18)
[…]
According to the Lagrange’s Mean Value Theorem, there must exist a number $c\in[a,x]$ with
$\frac{{\rm d^{n+1}}f}{{\rm d}c^{n+1}}=\frac{\int_{a}^{x}\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}\cdot(x-t)^{n}}{\frac{(x-a)^{n+1}}{n+1}},$ (19)
this gives us
$\int_{a}^{x}\frac{{\rm d^{n+1}}f}{{\rm d}t^{n+1}}\cdot(x-t)^{n}=\frac{{\rm d^{n+1}}f}{{\rm d}c^{n+1}}\cdot\frac{(x-a)^{n+1}}{n+1}$ (20)
[…]
and the equation on the left side of the proposition is proved completely. The inequality on the right-hand side is clearly established by the Lagrange’s Mean Value Theorem. $\square$
Lemma A.1 (Error Bound for the Midpoint Rule)
Suppose that $f(x)$ is a $m$-th $(m\geq 2)$ order differentiable function on the interval (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), and $K=\sup\left\{\left|\frac{\mathrm{d}^{2}f}{\mathrm{d}x^{2}}\right|:x\in[a,b]\right\}\in\mathbb{R}$, then
$\left|\int_{a}^{b}f(x)~{}\mathrm{d}x-(b-a)f\left(\frac{a+b}{2}\right)\right|\leq\frac{K}{24}(b-a)^{3}$
Proof.  We do the first order Taylor Expansion for $f(x)$ at the midpoint $x=\frac{a+b}{2}$ of the interval $[a,b]$:
$\left|\int_{a}^{b}f(x)~{}\mathrm{d}x-(b-a)f\left(\frac{a+b}{2}\right)\right|=\left|\int_{a}^{b}\left[f(x)-f\left(\frac{a+b}{2}\right)\right]~{}\mathrm{d}x\right|$ (21)
$=$ $\left|\int_{a}^{b}\left[\frac{\mathrm{d}f}{\mathrm{d}\left(\frac{a+b}{2}\right)}\cdot\left(x-\frac{a+b}{2}\right)+R_{1}(x)\right]\mathrm{d}x\right|$
$=$ $\left|0+\int_{a}^{b}R_{1}(x)~{}\mathrm{d}x\right|=\left|\int_{a}^{b}R_{1}(x)~{}\mathrm{d}x\right|$
Following Theorem A.1, we do the inequality scaling as below:
$\left|\int_{a}^{b}R_{1}(x)~{}\mathrm{d}x\right|\leq\left|\int_{a}^{b}\frac{K}{2!}\left(x-\frac{a+b}{2}\right)^{2}~{}\mathrm{d}x\right|=\frac{K}{6}\left[\left(\frac{b-a}{2}\right)^{3}-\left(\frac{a-b}{2}\right)^{3}\right]=\frac{K}{24}(b-a)^{3}$ (22)
$\square$
Lemma A.2 (Differential-Integral Error Order Estimation)
Suppose that $f(x)$ is a $m$-th $(m\geq 3)$ order differentiable function on the interval (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) and $K_{i}=\sup\left\{\left|\frac{\mathrm{d}^{i}f}{\mathrm{d}x^{i}}\right|:x\in(-\infty,+\infty)\right\}<+\infty~{}(i=1,2,\cdots,m)$, then on the closed interval $[a,b]$, we have
$\mathcal{A}:=\left\{\left|\sum_{i=1}^{L}\left|f\left(a+\frac{b-a}{L}i\right)-f\left(a+\frac{b-a}{L}(i-1)\right)\right|-\int_{a}^{b}\left|\frac{\mathrm{d}f}{\mathrm{d}x}\right|~{}\mathrm{d}x\right|\right\}\leq R(L)\sim o(\frac{1}{L}).$
Proof.  For each sub-interval $\left(a+\frac{b-a}{L}(i-1),a+\frac{b-a}{L}i\right)$, we perform a Taylor Expansion at the midpoint $x=a+\frac{b-a}{L}(i-\frac{1}{2})$ to obtain the following expression:
$\sum_{i=1}^{L}\left|f(a+\frac{b-a}{L}i)-f(a+\frac{b-a}{L}(i-1))\right|$ (23)
$=$ $\sum_{i=1}^{L}\left|\left[f\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)+\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{2L}(i-\frac{1}{2})\right)}\cdot\frac{b-a}{L}+\mathcal{O}\left[\left(\frac{b-a}{2L}\right)^{2}\right]\right]\right.$
$\left.-\left[f\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)+\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{2L}(i-\frac{1}{2})\right)}\cdot\frac{a-b}{L}+\mathcal{O}\left[\left(\frac{b-a}{2L}\right)^{2}\right]\right]\right|$
$=$ $\sum_{i=1}^{L}\left|\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)}\cdot\frac{b-a}{L}+\mathcal{O}(\frac{1}{L^{2}})\right|$
$\leq$ $\frac{b-a}{L}\cdot\sum_{i=1}^{L}\left|\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)}\right|+\sum_{i=1}^{L}\mathcal{O}(\frac{1}{L^{2}})$
$=$ $\frac{b-a}{L}\cdot\sum_{i=1}^{L}\left|\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)}\right|+\mathcal{O}(\frac{1}{L})$
Following Lemma A.1, we do the inequality scaling for the first-order derivative summation:
$\frac{b-a}{L}\cdot\sum_{i=1}^{L}\left|\frac{\mathrm{d}f}{\mathrm{d}\left(a+\frac{b-a}{L}(i-\frac{1}{2})\right)}\right|\leq\sum_{i=1}^{L}\int_{a+\frac{b-a}{L}(i-1)}^{a+\frac{b-a}{L}i}\left[\left|\frac{\mathrm{d}f}{\mathrm{d}x}\right|+\frac{K_{3}}{24}\left(\frac{b-a}{L}\right)^{3}\right]~{}\mathrm{d}x$ (24)
$=$ $\int_{a}^{b}\left[\left|\frac{\mathrm{d}f}{\mathrm{d}x}\right|+\frac{K_{3}(b-a)^{3}}{24L^{2}}\right]~{}\mathrm{d}x=\int_{a}^{b}\left|\frac{\mathrm{d}f}{\mathrm{d}x}\right|~{}\mathrm{d}x+\frac{K_{3}(b-a)^{4}}{24L^{2}}$
$\mathcal{A}:\leq\frac{K_{3}(b-a)^{4}}{24L^{2}}+\mathcal{O}(\frac{1}{L})=R(L)\sim o(\frac{1}{L})$ (25)
$\square$
A.3 Main Theorem Proof
Finally, we formally prove our main theorem A.1 in this sub-section, thereby verifying the Hypotheses 1 in the main paper. Since the $d$ components of ${\boldsymbol{V}}(s_{i})-{\boldsymbol{V}}(s_{j})$ are independent of each other, we only consider the $d$-th dimensional component and the rest of the components $V_{d}(s_{i})-V_{d}(s_{j})$ can be proved in the same way.
Proof.  According to Lemma A.2, we can approximate $V_{d}(s_{i})-V_{d}(s_{j})$ as
$V_{d}(s_{i})-V_{d}(s_{j})$ $=\sum_{l=1}^{L}\left|F_{id}(l)-F_{id}(l-1)\right|-\sum_{l=1}^{L}\left|F_{jd}(l)-F_{jd}(l-1)\right|$ (26)
$=\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x+\mathcal{O}(\frac{1}{L})$
$\approx\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x$
Now we want to remove the absolute value of the integrated function. Due to the continuity of the first-order derivative, there must be several zero points that are not extreme points, dividing the domain of function definition into several open intervals with alternating constant positive and negative function values. We first define such a set of zeros on the domain of definition $\mathcal{D}=[1,L]$:
$\mathcal{X}_{i}$ =$=\left\{x~{}|~{}x\in\mathcal{D};\frac{\mathrm{d}F_{id}}{\mathrm{d}x}=0\right\}=\{i_{1},i_{2},\cdots,i_{p}\},$ (27)
$\mathcal{X}_{j}$ =$=\left\{x~{}|~{}x\in\mathcal{D};\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}=0\right\}=\{j_{1},j_{2},\cdots,j_{q}\}.$
For all zero points $i_{t}\in\mathcal{X}_{i}$ or $j_{t}\in\mathcal{X}_{j}$, they must satisfy the following properties to ensure that they are not extreme points:
$\exists\epsilon>0,\forall{\Delta x<\epsilon},\frac{\mathrm{d}F_{id}}{\mathrm{d}(i_{t}-\Delta x)}\cdot\frac{\mathrm{d}F_{id}}{\mathrm{d}(i_{t}+\Delta x)}<0,$ (28)
and $j_{t}$ is the same as $i_{t}$. Therefore, in the sub-interval $(1,i_{1}),(i_{1},i_{2}),\cdots,(i_{p},L)$, $F_{id}(x)$ is alternately constant positive and constant negative, and the same as $F_{jd}(x)$.
We assume that the first interval $(1,i_{1})$ and $(1,j_{1})$ are constant positive intervals (the same can be proven for constant negative intervals). In this setting, we can continue to simplify Eq.(26):
$\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x$ (29)
$=$ $~{}\left(\int_{1}^{i_{1}}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x+\int_{i_{1}}^{i_{2}}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x+\cdots+\int_{i_{p}}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x\right)$
$-\left(\int_{1}^{j_{1}}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x+\int_{j_{1}}^{j_{2}}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x+\cdots+\int_{j_{q}}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x\right)$
$=$ $~{}\left[\left.F_{id}(x)\right|_{1}^{i_{1}}+\left.(-1)\cdot F_{id}(x)\right|_{i_{1}}^{i_{2}}+\cdots+(-1)^{i_{p}}\cdot\left.F_{id}(x)\right|_{i_{p}}^{L}\right]$
$-\left[\left.F_{jd}(x)\right|_{1}^{j_{1}}+\left.(-1)\cdot F_{jd}(x)\right|_{j_{1}}^{j_{2}}+\cdots+(-1)^{j_{q}}\cdot\left.F_{jd}(x)\right|_{j_{q}}^{L}\right]$
$=$ $~{}\left[-F_{id}(1)+F_{jd}(1)\right]+\left[2\sum_{k=1}^{p}(-1)^{k-1}\cdot F_{id}(i_{k})-2\sum_{k=1}^{q}(-1)^{k-1}\cdot F_{jd}(j_{k})\right]$
$+\left[(-1)^{i_{p}-1}\cdot F_{id}(L)-(-1)^{j_{q}-1}\cdot F_{jd}(L)\right]$
Since $F_{id}(x)$ and $F_{jd}(x)$ are taken from the functional space, $\{F_{id}(1),\{F_{id}(i_{k})\}_{k=1}^{p}\}$ and $\{F_{jd}(1),\{F_{jd}(j_{k})\}_{k=1}^{q}\}$ can be seen as independent variables. We let
$\boldsymbol{c}=\left[(-1)^{i_{p}-1}\cdot F_{id}(L)-(-1)^{j_{q}-1}\cdot F_{jd}(L)\right],$ (30)
then we can rewrite Eq.(29) to the matrix form:
$\int_{1}^{L}\left|\frac{\mathrm{d}F_{id}}{\mathrm{d}x}\right|~{}\mathrm{d}x-\int_{1}^{L}\left|\frac{\mathrm{d}F_{jd}}{\mathrm{d}x}\right|~{}\mathrm{d}x=\boldsymbol{A}\boldsymbol{x}+\boldsymbol{c},$ (31)
where $\boldsymbol{A}$ is the coefficient matrix, and $\boldsymbol{x}$ is the unknown variable vector:
$\boldsymbol{A}=[-1,1,\underbrace{2,-2,2,-2,...,(-1)^{p-1}\cdot 2}_{p~{}\text{dimensions}},\underbrace{-2,2,-2,2,...,(-1)^{q}\cdot 2}_{q~{}\text{dimensions}}]\in\mathbb{R}^{1\times(p+q+2)}$ (32)
$\boldsymbol{x}=[F_{id}(1),F_{jd}(1),F_{id}(i_{1}),...,F_{id}(i_{p}),F_{jd}(j_{1}),...,F_{jd}(j_{q})]^{\top}\in\mathbb{R}^{(p+q+2)\times 1}$
We let $i_{p}\equiv j_{q}(\mathrm{mod}~{}2)$, and return to the condition in the mathematical expectation on both sides of the inequality in the main theorem A.1 for the following categorical discussion:
  • If $F_{id}(L)=F_{jd}(L)=c_{d}$, then $\boldsymbol{c}=\boldsymbol{0}$, so $V_{d}(s_{i})-V_{d}(s_{j})=0\Longrightarrow\boldsymbol{Ax}=\boldsymbol{0}$;
[…]
We denote $N(A)$ as the solution space of $\{\boldsymbol{x}|\boldsymbol{Ax}=\boldsymbol{0}\}$ and $P(A)$ as the solution space of $\{\boldsymbol{x}|\boldsymbol{Ax}=-\boldsymbol{c}\}$. For $x\in P(A)$, $x=x_{p}$ or $x=x_{p}+x_{n}$ where $x_{p}$ is the special solution and $x_{n}$ is the solution in zero space, i.e., $x_{n}\in N(A)$. Since $\mathrm{rank}(A)=1<p+q+2$, the special solution $x_{p}$ exists, so the size of the solution space $P(A)$ is larger than $N(A)$. In this case, when the variable $\boldsymbol{x}$ is sampled from the real number space, its probability of being in $N(A)$ is smaller, namely its probability of not being in $N(A)$ is larger. This is exactly equivalent to the form of the proof of Main Theorem A.1, so the proof is complete. $\square$
A.4 Extended Conclusion
During the proof of the main theorem, we can unlock a hidden conclusion due to the embedding interpolation: GLMs with a larger number of hidden layers may achieve more stable detection performance, i.e., discrepancies in embedding volatility of different samples will be more obvious.
In Lemma A.2, we have proved the upper bound of differential-integral error order estimation is the equivalent infinitesimals of $1/L$. Actually, when $L\rightarrow+\infty$, we have
$\lim_{L\rightarrow+\infty}\sum_{i=1}^{L}\left|f\left(a+\frac{b-a}{L}i\right)-f\left(a+\frac{b-a}{L}(i-1)\right)\right|=\int_{a}^{b}\left|\frac{\mathrm{d}f}{\mathrm{d}x}\right|~{}\mathrm{d}x,$ (33)
which is clear according to the definition of Riemann Sum. This means that when the value of $L$ increases, the differential-integral approximation error will be reduced, so the conclusion of the main theorem will be more accurate. In our experiments, we use GPT2-XL (48 layers) and Llama2-7B (32 layers) as training backbones and find that the average performance of GPT2-XL is higher and more stable than that of Llama2-7B, while the number of layers of GPT2-XL is 1.5 times that of Llama2-7B. This further validates the correctness of our extended conclusions.
Appendix B Motivation II: Empirical Analysis
B.1 Visualization Projection Setting
We use UMAP [38] for projection visualization, which is a nonlinear dimensionality reduction technique for mapping high-dimensional data into low-dimensional spaces. It uses optimization of the local and global structure of the original data to produce a high-quality mapping that can preserve the original data’s local and global characteristics.
The UMAP algorithm consists of two steps: (i) Calculate the local similarity between each data point and its immediate neighbors to construct a local similarity map. (ii) Map the high-dimensional data to a low-dimensional space by optimizing an objective function that maintains the local and global structure. The hyperparameter “n_neighbors” in the first step is key, we set it as 10 in our paper.
B.2 Visualization Example
We select the MATH [7] dataset for mathematical reasoning and the OPUS [52] for text generation (translation). In the MATH, we choose four domains: algebra, geometry, number theory, and precalculus; In the OPUS, we also choose four domains: Bible, news, TED, and health. We present examples of inputs and outputs under different domains for mathematical reasoning and translation scenarios in Table 4 and Table 5, respectively. They correspond to the case projection of Figure 1. Obviously, under mathematical reasoning, the outputs under different domains may appear exactly the same, while it is impossible under translation.
Input Output
Domain 1: Algebra
Suppose that $f$ is a function and $f^{-1}$ is the inverse of $f$. If $f(1)=2$, $f(2)=6$, and $f(3)=5$, then what is $f^{-1}(f^{-1}(6))$? 1
Domain 2: Geometry
Two chords, $AB$ and $CD,$ meet inside a circle at $P.$ If $AP=CP=7,$ then what is $\frac{BP}{DP}$? 1
Domain 3: Number Theory
The product of the proper positive integer factors of $n$ can be written as $n^{(ax+b)/c}$, where $x$ is the number of positive divisors $n$ has, $c$ is a positive integer, and the greatest common factor of the three integers $a$, $b$, and $c$ is $1$. What is $a+b+c$? 1
Domain 4: Precalculus
Simplify $\frac{1}{1-\tan^{2}x}+\frac{1}{1-\cot^{2}x}.$ 1
Table 4: Examples of inputs/outputs from different domains in the mathematical reasoning scenario.
Input Output
Domain 1: Bible
The earth yielded grass, herbs yielding seed after their kind, and trees bearing fruit, with its seed in it, after their kind; and God saw that it was good. 于是地发生了青草,和结种子的菜蔬,各从其类;并结果子的树木,各从其类;果子都包着核。神看着是好的。
Domain 2: News
Supporters say the tunnels would benefit the environment and offer Californians a more secure water supply. 支持者们表示,这两条隧道将使环境受益并帮助加利福尼亚州确保供水更加安全。
Domain 3: TED
(Applause) June Cohen: Frank, that was beautiful, so touching. (鼓掌) 主持:Frank 刚刚真是太美丽、太感人啦!
Domain 4: Health
The White House Coronavirus Task Force was established on January 29. 白宫冠状病毒工作组(White House Coronavirus Task Force)是成立于2020年1月29日的以应对美国COVID-19疫情的工作组。
Table 5: Examples of inputs/outputs from different domains in the translation scenario.
Appendix C Algorithm: Pseudo-Code of TV Score Computation
The pseudo-code of our TV score computation pipeline is shown in Algorithm 1.
Algorithm 1 Trajectory Volatility (TV) Score Computation
0:  $L$: The number of hidden layers $N$: The size of ID dataset $k$: the smoothing differential order of TV score $\{{\boldsymbol{y}_{l}}\}_{1\leq l\leq L}$: the average hidden embedding of the OOD sample in each layer $\{[\hat{{\boldsymbol{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$: the average hidden embedding of all $N$ ID samples in each layer
1:  for $l\leftarrow 1$ to $L$ do
2:     Fitting ID samples $\{[\hat{{\boldsymbol{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$ to Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l},{\boldsymbol{\Sigma}}_{l})$
3:     $f({\boldsymbol{y}_{l}})\leftarrow\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)$
4:  end for
5:  for $i\rightarrow 1$ to $k$ do
6:     for $l\leftarrow 1$ to $L$ do
7:        $\Delta^{(i)}\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l}^{(i)},{\boldsymbol{\Sigma}}_{l}^{(i)})\leftarrow\mathcal{N}(\sum_{t=0}^{i}(-1)^{i+t}\mathrm{C}_{i}^{t}{\boldsymbol{\mu}}_{l+i},~{}\sum_{t=0}^{i}\mathrm{C}_{i}^{t}{\boldsymbol{\Sigma}}_{l+i})$
8:     end for
9:     $f^{(i)}({\boldsymbol{y}_{l}})\leftarrow\left({\boldsymbol{y}_{l}}^{(i)}-{\boldsymbol{\mu}_{l}}^{(i)}\right)^{\top}\left(\Sigma_{l}^{(i)}\right)^{-1}\left({\boldsymbol{y}_{l}}^{(i)}-{\boldsymbol{\mu}_{l}}^{(i)}\right)$
10:  end for
11:  for $i\rightarrow 1$ to $n$ do
12:     $\textrm{TV}_{i}\leftarrow\texttt{average}\left[f^{(i)}({\boldsymbol{y}_{l}})\right]_{1\leq l\leq L}$
13:  end for
13:  $\{\textrm{TV}_{i}\}_{1\leq i\leq k}$
Appendix D Experimental Setting Details
D.1 Basic Information of Dataset
For the ID dataset, we use the MultiArith [47], which consists of Math Word Problems on arithmetic reasoning. For the OOD datasets, we intuitively introduce two types of detection scenarios following [46]: (i) Far-shift OOD setting, we select the MATH [13] as the OOD data, which spans across distinct mathematical domains encompassing algebra, geometry, counting and probability, number theory, and precalculus. It contains college difficulty level math problems, whereas MultiArith has only elementary school difficulty, and thus can be considered as sourced from far-different distributions; (ii) Near-shift OOD setting, we select five arithmetic reasoning datasets as the OOD data: GSM8K [7], SVAMP [42], AddSub [15], SingleEq [21], and SingleOp [21], they all consist of Math Word Problems like the MultiArith but require different reasoning hops and knowledge points for solving the problems, and thus can be considered as sourced from near-different distributions. In addition, we present the data sizes and examples of all ID and OOD datasets, as shown in Table 6.
In-Distribution Dataset
MultiArith (Data Size: 600) Q: Kaleb was collecting cans for recycling. On Saturday he filled 5 bags up and on Sunday he filled 5 more bags. If each bag had 4 cans in it, how many cans did he pick up total? A: 40.0
Far Shift Out-of-Distribution Dataset Near Shift Out-of-Distribution Dataset
MATH-Algebra (Data Size: 1187) Q: How many real numbers are not in the domain of the function f(x) = 1x-64 + 1x2-64 + 1x3-64 ? A: 4 GSM8K (Data Size: 1318) Q: Judy teaches 5 dance classes, every day, on the weekdays and 8 classes on Saturday. If each class has 15 students and she charges \$15.00 per student, how much money does she make in 1 week? A: 7425
MATH-Geometry (Data Size: 479) Q: Suppose we are given seven points that are equally spaced around a circle. If $P$, $Q$, and $R$ are chosen to be any three of these points, then how many different possible values are there for $m\angle PQR$? A: 5 SVAMP (Data Size: 1000) Q: A mailman has to give 38 pieces of junk mail to each of the 78 blocks. If there are 19 houses on a block. How many pieces of junk mail should he give each house? A: 2.0
MATH-Counting and Probability (Data Size: 474) Q: Amy’s grandmother gave her 3 identical chocolate chip cookies and 4 identical sugar cookies. In how many different orders can Amy eat the cookies such that either she eats a chocolate chip cookie first, she eats a chocolate chip cookie last, or both? A: 25 AddSub (Data Size: 395) Q: While taking inventory at her pastry shop, Kelly realizes that she had 0.4 box of baking powder yesterday, but the supply is now down to 0.3 box. How much more baking powder did Kelly have yesterday?
A: 0.1
MATH-Number Theory (Data Size: 540) Q: Notice that $31\cdot 37=1147$ Find some integer $n$ with $22930\leq n<$ 2293 such that $31n\equiv 3\pmod{2293}.$ A: 222 SingleEq (Data Size: 508) Q: Fred had 7 dimes in his bank. His sister borrowed 3 of his dimes. How many dimes does Fred have now?
A: 4.0
MATH-Precalculus (Data Size: 546) Q: Let $\mathbf{a}=\begin{pmatrix}1\\ 3\\ 0\end{pmatrix},$ $\mathbf{b}=\begin{pmatrix}-11\\ 5\\ 2\end{pmatrix},$ and $\mathbf{c}=\begin{pmatrix}1+\sqrt{5}\\ 4\\ -5\end{pmatrix}.$ Find $k$ if the vectors $\mathbf{a}+\mathbf{b}+\mathbf{c}$ and $3(\mathbf{b}\times\mathbf{c})-8(\mathbf{c}\times\mathbf{a})+k(\mathbf{a}\times\mathbf{b})$ are orthogonal. A: 5 SingleOp (Data Size: 562) Q: Pamela starts with 30 bottle caps. Jean takes 26 away. How many bottle caps does Pamela end with?
A: 4.0
Table 6: ID and OOD datasets used in this paper.
D.2 ID Dataset Split
For the ID dataset MultiArith, we find that every 100 consecutive samples show a similar quadratic operation pattern (e.g., samples with id 0-100 are a mixture of addition and subtraction, and id 100-200 are a mixture of addition and multiplication). Therefore, we divide it into 6 subsets (6*100) in order. In each subset, we take the first 60 as training samples and the last 40 as test samples.
D.3 Training Implementation
We train Llama2-7B [53] and GPT2-XL [6] models on the training split of MultiArith. Llama2-7B is trained with AdamW optimizer [33] for 10K steps and 8 batch sizes in 4-card RTX 3090 (2 per card). The learning rate is set to 1e-5, the warmup step to 10, and the maximum gradient normalization to 0.3. GPT2-XL is trained for 3K steps and 128 batch sizes in a single RTX 3090, and other configurations are the same as Llama2-7B.
D.4 OOD Dataset Selection Rationality
In this part, we examine the rationality of the OOD data selection, ensuring that the OOD data distribution significantly differs from the pre-trained data distribution and has not been fully learned during the pre-training phase. Some research [57, 59] have confirmed the absence of data leakage in Llama2 for MATH and GSM8K datasets, we still conduct experiments and analyses to ensure this.
For GPT2-XL, We can determine this from the time dimension. GPT-2 was released in 2019, but the MATH and GSM8k datasets were released in 2021, so they are unlikely to appear in the pre-training data. However, for Llama2-7B, due to the closed-source data, we cannot confirm which data the model used in the pre-training phase, so we cannot fully determine whether the selected dataset was OOD for the model from a data perspective.
Therefore, we argue that a dataset can be considered OOD when it is beyond the capability of the base model as claimed by prior work [50, 31]. We test all ten datasets we select as the OOD data in the pre-trained Llama2-7B and GPT2-XL model, and Table 7 shows the results. We find that GPT2-XL cannot handle any of the ten mathematical reasoning tasks, and Llama2-7B performs with very low accuracy. Therefore, from a capability standpoint, we can ensure that these datasets are OOD for these two GLMs.
Far-shift OOD Setting Near-shift OOD Setting
Dataset Llama2-7B GPT2-XL Dataset Llama2-7B GPT2-XL
Accuracy of pre-trained model Accuracy of pre-trained model
Algebra 6 / 1187 0 / 1187 GSM8K 0 / 1318 0 / 1318
Geometry 2 / 479 0 / 479 SVAMP 8 / 1000 0 / 1000
Cnt.&Prob 4 / 474 0 / 474 AddSub 13 / 395 0 / 1000
Num.Theory 0 / 540 0 / 540 SingleEq 5 / 395 0 / 508
Precalculus 1 / 546 0 / 540 SingleOp 7 / 508 0 / 508
Table 7: Accuracies of all datasets we select as the OOD data in pre-trained GLMs.
D.5 Baseline
Let ${\boldsymbol{x}}$ represent the input sequence and ${\boldsymbol{y}}$ the output sequence, with lengths denoted as $n_{x}$ and $n_{y}$ respectively. In addition, we assume that $p(\cdot;{\boldsymbol{\theta}})$ represents the GLM parameterized by ${\boldsymbol{\theta}}$ that has been trained in ID dataset $\mathcal{D}$, outputting a sequence of softmax probabilities. We compare some training-free baseline methods as below:
  • Maximum Softmax Probability [14]:
    $\sum_{i=1}^{n_{y}}\max\frac{p(y_{i}|y_{\prec i},{\boldsymbol{x}};{\boldsymbol{\theta}})}{n_{y}}.$ (34)
  • Monte-Carlo Dropout [10]:
    $\int p({\boldsymbol{y}}|{\boldsymbol{x}};{\boldsymbol{\theta}})q({\boldsymbol{\theta}})\mathrm{d}\theta,$ (35)
    where $q({\boldsymbol{\theta}})\approx p({\boldsymbol{\theta}}|\mathcal{D})$.
  • Sequence Perplexity [46]:
    $\left[\prod_{i=1}^{n_{y}}p(y_{i}|y_{\prec i},{\boldsymbol{x}};{\boldsymbol{\theta}})\right]^{-\frac{1}{N}}.$ (36)
  • Input Embedding [46]:
    $({\boldsymbol{x}}-{\boldsymbol{\mu}_{x}})^{\top}{\boldsymbol{\Sigma}_{x}^{-1}}({\boldsymbol{x}}-{\boldsymbol{\mu}_{x}}),$ (37)
    where ${\boldsymbol{\mu}_{x}}$ and ${\boldsymbol{\Sigma}_{x}}$ represent the mean and variance of the Gaussian distribution associated with the first-layer hidden state.
  • Output Embedding [46]:
    $({\boldsymbol{y}}-{\boldsymbol{\mu}_{y}})^{\top}{\boldsymbol{\Sigma}_{y}^{-1}}({\boldsymbol{y}}-{\boldsymbol{\mu}_{y}}),$ (38)
    where ${\boldsymbol{\mu}_{y}}$ and ${\boldsymbol{\Sigma}_{y}}$ represent the mean and variance of the Gaussian distribution associated with the final-layer hidden state.
Appendix E Full Main Experimental Results
OOD Detection.
AUROC and FPR95 results of each dataset are shown in Table 8 (Llama2-7B) and Table 9 (GPT2-XL).
OOD Quality Estimation.
Kendall’s $\tau$ correlation results of each dataset are shown in Table 10 (Llama2-7B) and Table 11 (GPT2-XL).
Far-shift OOD Setting
Method Dataset Algebra Geometry Counting and Probability Number Theory Precalculus Average
AUROC $\uparrow$  /   FPR95 $\downarrow$
MS-Prob [14] 79.97$\pm$1.60 / 80.97$\pm$4.90 82.60$\pm$1.25 / 82.57$\pm$4.62 63.87$\pm$1.64 / 96.19$\pm$1.57 76.20$\pm$1.46 / 85.20$\pm$2.69 90.68$\pm$0.94 / 62.27$\pm$4.03 78.66$\pm$1.38 / 81.44$\pm$3.56
MC-Drop [10] 72.12$\pm$1.68 / 85.43$\pm$6.60 74.23$\pm$1.74 / 81.17$\pm$6.28 55.64$\pm$2.87 / 97.15$\pm$1.34 67.53$\pm$2.08 / 89.09$\pm$5.41 73.61$\pm$2.66 / 82.36$\pm$4.78 68.63$\pm$2.21 / 87.04$\pm$4.88
PPL [3] 84.17$\pm$1.43 / 52.75$\pm$5.13 88.67$\pm$1.36 / 53.32$\pm$5.38 77.42$\pm$1.97 / 71.28$\pm$3.90 83.88$\pm$2.13 / 68.90$\pm$4.22 94.05$\pm$0.43 / 19.03$\pm$3.17 85.64$\pm$1.46 / 53.06$\pm$4.36
I-Emb [46] 81.51$\pm$1.09 / 62.84$\pm$4.56 75.41$\pm$0.97 / 69.66$\pm$2.07 62.53$\pm$1.30 / 88.35$\pm$1.72 84.42$\pm$0.85 / 54.80$\pm$7.25 75.59$\pm$0.92 / 63.71$\pm$2.87 75.89$\pm$1.03 / 67.87$\pm$3.69
O-Emb [46] 76.95$\pm$1.54 / 74.92$\pm$3.02 78.72$\pm$1.31 / 80.97$\pm$2.23 61.43$\pm$1.40 / 88.45$\pm$1.58 70.23$\pm$1.36 / 80.18$\pm$1.71 86.97$\pm$1.32 / 51.51$\pm$2.27 74.86$\pm$1.39 / 75.21$\pm$2.16
TV score (Ours) 98.87$\pm$0.16 / 4.67$\pm$1.41 99.03$\pm$0.09 / 3.70$\pm$0.42 97.70$\pm$0.15 / 8.83$\pm$1.33 98.43$\pm$0.13 / 7.37$\pm$1.46 99.78$\pm$0.02 / 1.47$\pm$0.20 98.76$\pm$0.11 / 5.21$\pm$0.98
w/ DiSmo (Ours) 94.71$\pm$0.93 / 39.65$\pm$7.28 94.08$\pm$0.80 / 50.52$\pm$6.14 83.08$\pm$1.28 / 80.07$\pm$1.57 94.57$\pm$0.75 / 37.74$\pm$7.58 99.82$\pm$0.02 / 1.11$\pm$0.19 93.25$\pm$0.76 / 41.82$\pm$4.69
$\Delta$ (bold - underline) +14.70 / -48.08 +10.36 / -49.62 +20.28 / -62.45 +14.01 / -47.43 +5.77 / -17.92 +13.12 / -47.85
Near-shift OOD Setting
Method Dataset GSM8K SVAMP AddSub SingleEq SingleOp Average
AUROC $\uparrow$  /   FPR95 $\downarrow$
MS-Prob [14] 53.08$\pm$1.67 / 94.07$\pm$1.86 56.56$\pm$1.53 / 90.06$\pm$2.39 63.31$\pm$1.88 / 87.36$\pm$2.42 66.68$\pm$1.32 / 86.15$\pm$2.61 61.07$\pm$1.30 / 86.91$\pm$2.76 60.14$\pm$1.54 / 88.91$\pm$2.41
MC-Drop [10] 48.87$\pm$2.43 / 96.78$\pm$1.21 44.90$\pm$2.76 / 92.33$\pm$2.09 57.34$\pm$1.57 / 89.15$\pm$2.34 54.09$\pm$2.42 / 90.07$\pm$1.96 56.46$\pm$1.85 / 91.21$\pm$1.83 52.33$\pm$2.21 / 91.92$\pm$1.89
PPL [3] 52.24$\pm$2.57 / 95.56$\pm$1.21 55.12$\pm$1.80 / 89.24$\pm$2.09 62.88$\pm$1.76 / 80.96$\pm$2.34 67.14$\pm$1.34 / 81.38$\pm$1.96 59.39$\pm$1.98 / 83.30$\pm$1.83 59.35$\pm$1.89 / 86.09$\pm$1.89
I-Emb [46] 45.68$\pm$1.50 / 95.05$\pm$1.17 60.92$\pm$1.34 / 86.97$\pm$4.13 66.28$\pm$0.92 / 76.09$\pm$2.93 61.18$\pm$1.22 / 87.36$\pm$1.89 67.60$\pm$1.20 / 77.78$\pm$2.53 60.33$\pm$1.37 / 84.65$\pm$2.53
O-Emb [46] 35.39$\pm$1.24 / 91.22$\pm$1.27 36.77$\pm$1.14 / 90.41$\pm$1.61 63.08$\pm$0.92 / 77.12$\pm$3.11 43.70$\pm$1.02 / 86.69$\pm$0.91 43.55$\pm$0.99 / 86.87$\pm$1.04 44.50$\pm$1.06 / 86.46$\pm$1.59
TV score (Ours) 94.88$\pm$0.25 / 14.22$\pm$1.64 94.51$\pm$0.20 / 12.89$\pm$1.11 85.84$\pm$1.06 / 82.63$\pm$2.22 93.97$\pm$0.24 / 17.39$\pm$1.09 94.00$\pm$0.20 / 14.61$\pm$0.85 92.64$\pm$0.39 / 28.39$\pm$1.38
w/ DiSmo (Ours) 55.21$\pm$1.91 / 95.71$\pm$0.88 38.24$\pm$1.53 / 87.06$\pm$0.94 87.06$\pm$0.94 / 71.02$\pm$4.33 56.66$\pm$1.34 / 93.76$\pm$1.28 47.46$\pm$1.32 / 92.48$\pm$1.13 56.99$\pm$1.41 / 88.01$\pm$1.71
$\Delta$ (bold - underline) +41.80 / -77.00 +33.59 / -74.08 +20.78 / -5.07 +26.83 / -63.99 +26.40 / -63.17 +32.31 / -56.26
Table 8: OOD Detection — Offline Discrimination (Llama2-7B): AUROC and FPR95 results ($p$-value > $0.05$ are grayed out). Underline denote SOTA among all baselines and bold denote SOTA among all methods.
Far-shift OOD Setting
Method Dataset Algebra Geometry Counting and Probability Number Theory Precalculus Average
AUROC $\uparrow$  /   FPR95 $\downarrow$
MS-Prob 72.13$\pm$1.42 / 78.35$\pm$2.35 64.09$\pm$1.78 / 87.42$\pm$1.24 68.45$\pm$1.56 / 82.35$\pm$1.57 69.37$\pm$1.43 / 82.09$\pm$1.90 78.67$\pm$0.89 / 61.23$\pm$3.02 70.54$\pm$1.42 / 78.29$\pm$2.02
MC-Drop 62.41$\pm$2.10 / 88.02$\pm$1.79 66.63$\pm$1.82 / 82.49$\pm$1.12 59.32$\pm$2.03 / 92.34$\pm$1.54 63.88$\pm$1.92 / 86.71$\pm$1.95 70.19$\pm$1.47 / 73.90$\pm$1.87 66.18$\pm$1.87 / 84.69$\pm$1.65
PPL 84.24$\pm$1.01 / 63.78$\pm$2.38 82.17$\pm$0.89 / 67.90$\pm$2.45 79.12$\pm$1.20 / 72.03$\pm$1.23 72.40$\pm$1.37 / 71.65$\pm$1.45 86.19$\pm$0.75 / 47.27$\pm$2.98 80.82$\pm$1.04 / 64.53$\pm$2.10
I-Emb 86.23$\pm$0.78 / 46.12$\pm$1.45 83.20$\pm$0.94 / 53.29$\pm$2.06 79.58$\pm$1.43 / 60.45$\pm$2.95 89.44$\pm$0.64 / 52.49$\pm$1.87 92.86$\pm$0.39 / 34.32$\pm$2.16 86.26$\pm$0.84 / 49.33$\pm$2.10
O-Emb 78.13$\pm$1.14 / 64.46$\pm$2.78 77.98$\pm$1.17 / 68.82$\pm$3.26 71.07$\pm$1.67 / 84.63$\pm$4.24 80.25$\pm$0.97 / 56.23$\pm$4.05 82.34$\pm$0.86 / 54.08$\pm$2.79 77.95$\pm$1.16 / 65.64$\pm$3.42
TV score (Ours) 98.35$\pm$0.07 / 6.24$\pm$0.24 97.27$\pm$0.11 / 9.23$\pm$0.35 85.52$\pm$0.12 / 52.51$\pm$1.78 92.96$\pm$0.06 / 29.86$\pm$1.26 93.24$\pm$0.04 / 22.68$\pm$1.14 93.47$\pm$0.08 / 24.10$\pm$0.95
w/ DiSmo (Ours) 93.68$\pm$0.19 / 23.24$\pm$1.45 94.39$\pm$0.15 / 12.17$\pm$0.79 95.83$\pm$0.13 / 9.86$\pm$0.46 99.17$\pm$0.04 / 2.42$\pm$0.22 99.62$\pm$0.02 / 1.75$\pm$0.13 96.54$\pm$0.11 / 9.89$\pm$0.61
$\Delta$ (bold - underline) +12.12 / -39.88 +14.07 / -44.06 +16.25 / -50.59 +9.27 / -50.07 +6.24 / -32.57 +10.28 / -39.44
Near-shift OOD Setting
Method Dataset GSM8K SVAMP AddSub SingleEq SingleOp Average
AUROC $\uparrow$  /   FPR95 $\downarrow$
MS-Prob 54.06$\pm$1.58 / 92.45$\pm$2.37 65.50$\pm$1.43 / 73.62$\pm$4.18 70.60$\pm$0.92 / 77.56$\pm$2.76 79.96$\pm$0.87 / 57.38$\pm$1.95 65.47$\pm$1.18 / 80.37$\pm$2.02 67.12$\pm$1.20 / 76.27$\pm$2.66
MC-Drop 63.44$\pm$2.08 / 79.85$\pm$3.35 62.13$\pm$1.65 / 77.28$\pm$3.21 67.97$\pm$1.62 / 71.42$\pm$2.88 71.29$\pm$1.43 / 69.22$\pm$2.41 52.89$\pm$1.80 / 92.78$\pm$0.67 63.54$\pm$1.72 / 78.08$\pm$2.50
PPL 72.89$\pm$1.56 / 74.63$\pm$1.52 70.79$\pm$1.38 / 79.65$\pm$1.21 67.50$\pm$1.04 / 87.76$\pm$0.92 87.14$\pm$0.57 / 43.09$\pm$1.23 70.40$\pm$1.06 / 76.80$\pm$1.45 73.74$\pm$1.12 / 72.39$\pm$1.27
I-Emb 87.58$\pm$1.14 / 48.45$\pm$2.06 84.81$\pm$1.20 / 62.75$\pm$5.68 80.34$\pm$0.64 / 46.62$\pm$2.17 81.44$\pm$0.49 / 49.05$\pm$3.24 81.91$\pm$0.93 / 57.65$\pm$2.67 83.22$\pm$0.88 / 52.90$\pm$3.16
O-Emb 82.05$\pm$1.36 / 60.44$\pm$3.96 81.42$\pm$2.02 / 68.43$\pm$4.38 72.98$\pm$0.85 / 66.72$\pm$1.77 80.00$\pm$1.02 / 70.23$\pm$1.87 79.94$\pm$0.94 / 57.68$\pm$1.60 79.28$\pm$1.24 / 64.70$\pm$2.72
TV score (Ours) 98.26$\pm$0.06 / 1.78$\pm$0.04 98.99$\pm$0.02 / 1.13$\pm$0.03 80.56$\pm$0.82 / 52.23$\pm$1.29 97.35$\pm$0.23 / 13.91$\pm$0.45 99.12$\pm$0.02 / 0.07$\pm$0.01 94.86$\pm$0.23 / 13.82$\pm$0.36
w/ DiSmo (Ours) 97.62$\pm$0.21 / 3.58$\pm$0.16 92.19$\pm$0.44 / 13.07$\pm$1.68 83.35$\pm$0.57 / 41.70$\pm$1.29 97.83$\pm$0.03 / 9.94$\pm$0.32 99.94$\pm$0.02 / 0.03$\pm$0.01 94.19$\pm$0.25 / 13.66$\pm$0.69
$\Delta$ (bold - underline) +10.68 / -46.67 +14.18 / -61.62 +3.01 / -4.92 +10.69 / -33.15 +18.03 / -57.62 +11.64 / -39.24
Table 9: OOD Detection — Offline Discrimination (GPT2-XL): AUROC and FPR95 results ($p$-value > $0.05$ are grayed out). Underline denote SOTA among all baselines and bold denote SOTA among all methods.
Method Dataset ID + Far-shift OOD ID + Near-shift OOD
Algebra Geometry Cnt.&Prob. Num.Theory Precalculus Average GSM8K SVAMP AddSub SingleEq SingleOp Average
Kendall Rank Correlation Coefficient
MS-Prob 0.035$\pm$0.025 0.024$\pm$0.020 0.084$\pm$0.022 -0.043$\pm$0.021 0.021$\pm$0.014 0.024$\pm$0.020 0.064$\pm$0.022 0.067$\pm$0.015 0.076$\pm$0.016 0.057$\pm$0.019 -0.075$\pm$0.017 0.038$\pm$0.018
PPL -0.027$\pm$0.020 0.071$\pm$0.013 0.112$\pm$0.015 0.035$\pm$0.016 0.036$\pm$0.011 0.050$\pm$0.015 0.073$\pm$0.019 0.027$\pm$0.017 0.082$\pm$0.015 0.091$\pm$0.019 0.079$\pm$0.017 0.074$\pm$0.017
I-Emb 0.074$\pm$0.020 0.119$\pm$0.015 0.054$\pm$0.018 0.111$\pm$0.016 0.034$\pm$0.013 0.078$\pm$0.016 0.042$\pm$0.020 -0.025$\pm$0.016 0.052$\pm$0.014 0.055$\pm$0.020 0.058$\pm$0.018 0.036$\pm$0.018
O-Emb 0.050$\pm$0.023 0.078$\pm$0.019 0.064$\pm$0.020 0.065$\pm$0.020 0.034$\pm$0.009 0.058$\pm$0.018 0.042$\pm$0.018 -0.001$\pm$0.013 0.056$\pm$0.015 0.032$\pm$0.015 0.061$\pm$0.015 0.038$\pm$0.015
TV score (Ours) 0.182$\pm$0.016 0.116$\pm$0.011 0.191$\pm$0.011 0.174$\pm$0.014 0.142$\pm$0.009 0.161$\pm$0.012 0.146$\pm$0.010 0.209$\pm$0.019 0.195$\pm$0.018 0.145$\pm$0.017 0.101$\pm$0.020 0.159$\pm$0.017
w/ DiSmo (Ours) 0.095$\pm$0.020 0.059$\pm$0.014 0.123$\pm$0.020 0.131$\pm$0.016 0.148$\pm$0.009 0.111$\pm$0.016 0.178$\pm$0.025 0.113$\pm$0.015 0.121$\pm$0.016 0.079$\pm$0.017 0.076$\pm$0.019 0.113$\pm$0.018
$\Delta$ (bold - underline) +0.021 -0.003 +0.079 +0.109 +0.112 +0.083 +0.073 +0.142 +0.039 +0.055 +0.022 +0.085
Spearman Rank Correlation Coefficient
MS-Prob 0.027$\pm$0.021 0.057$\pm$0.019 0.039$\pm$0.021 0.067$\pm$0.021 0.004$\pm$0.017 0.038$\pm$0.020 0.071$\pm$0.021 -0.086$\pm$0.018 0.063$\pm$0.019 0.081$\pm$0.015 0.002$\pm$0.019 0.026$\pm$0.018
PPL 0.039$\pm$0.014 0.018$\pm$0.018 0.073$\pm$0.017 0.051$\pm$0.015 0.045$\pm$0.015 0.045$\pm$0.016 0.086$\pm$0.019 0.090$\pm$0.019 0.016$\pm$0.019 0.039$\pm$0.015 0.017$\pm$0.016 0.050$\pm$0.018
I-Emb 0.150$\pm$0.021 0.136$\pm$0.018 0.061$\pm$0.015 0.038$\pm$0.016 0.127$\pm$0.016 0.102$\pm$0.017 0.042$\pm$0.022 0.128$\pm$0.016 0.154$\pm$0.016 0.110$\pm$0.017 0.140$\pm$0.015 0.115$\pm$0.017
O-Emb 0.002$\pm$0.018 0.001$\pm$0.020 0.015$\pm$0.015 0.033$\pm$0.016 0.012$\pm$0.018 0.025$\pm$0.017 -0.040$\pm$0.017 -0.019$\pm$0.016 0.045$\pm$0.019 0.101$\pm$0.018 -0.027$\pm$0.017 0.012$\pm$0.017
TV score (Ours) 0.122$\pm$0.012 0.169$\pm$0.015 0.102$\pm$0.016 0.126$\pm$0.016 0.216$\pm$0.014 0.147$\pm$0.015 0.124$\pm$0.016 0.188$\pm$0.017 0.127$\pm$0.017 0.165$\pm$0.017 0.188$\pm$0.019 0.158$\pm$0.017
w/ DiSmo (Ours) 0.139$\pm$0.014 0.148$\pm$0.013 0.176$\pm$0.017 0.178$\pm$0.017 0.121$\pm$0.016 0.152$\pm$0.015 0.071$\pm$0.018 0.150$\pm$0.020 0.167$\pm$0.014 0.131$\pm$0.017 0.153$\pm$0.017 0.134$\pm$0.017
$\Delta$ (bold - underline) -0.011 +0.033 +0.103 +0.127 +0.089 +0.050 +0.038 +0.060 +0.013 +0.055 +0.048 +0.043
Table 10: OOD Quality Estimation (Llama2-7B): Kendall’s $\tau$ and Spearman correlation between various OOD scores and benchmark quality metric binary matching. Each column shows the correlation when ID and OOD samples are merged. Underline denotes the SOTA among all baselines, and bold denotes the SOTA among our methods.
Method Dataset ID + Far-shift OOD ID + Near-shift OOD
Algebra Geometry Cnt.&Prob. Num.Theory Precalculus Average GSM8K SVAMP AddSub SingleEq SingleOp Average
Kendall Rank Correlation Coefficient
MS-Prob 0.078$\pm$0.018 0.032$\pm$0.015 0.093$\pm$0.014 0.059$\pm$0.017 0.068$\pm$0.012 0.066$\pm$0.015 0.032$\pm$0.024 0.061$\pm$0.020 0.088$\pm$0.018 0.023$\pm$0.017 0.081$\pm$0.019 0.057$\pm$0.018
PPL 0.023$\pm$0.014 0.086$\pm$0.015 0.043$\pm$0.016 -0.009$\pm$0.015 0.036$\pm$0.012 0.036$\pm$0.014 0.068$\pm$0.022 0.008$\pm$0.022 0.035$\pm$0.018 0.001$\pm$0.015 0.064$\pm$0.015 0.035$\pm$0.018
I-Emb 0.084$\pm$0.012 0.052$\pm$0.011 0.089$\pm$0.013 0.062$\pm$0.011 0.007$\pm$0.010 0.059$\pm$0.012 0.041$\pm$0.019 0.010$\pm$0.020 -0.008$\pm$0.016 0.019$\pm$0.015 0.000$\pm$0.018 0.012$\pm$0.018
O-Emb 0.089$\pm$0.012 0.023$\pm$0.012 0.058$\pm$0.013 0.014$\pm$0.012 0.067$\pm$0.011 0.050$\pm$0.012 0.068$\pm$0.018 0.036$\pm$0.020 0.033$\pm$0.018 0.022$\pm$0.014 0.021$\pm$0.017 0.036$\pm$0.017
TV score (Ours) 0.127$\pm$0.010 0.142$\pm$0.010 0.158$\pm$0.009 0.098$\pm$0.011 0.167$\pm$0.009 0.138$\pm$0.010 0.087$\pm$0.013 0.166$\pm$0.019 0.110$\pm$0.015 0.152$\pm$0.013 0.141$\pm$0.016 0.131$\pm$0.015
w/ DiSmo (Ours) 0.134$\pm$0.010 0.112$\pm$0.009 0.132$\pm$0.010 0.161$\pm$0.009 0.158$\pm$0.009 0.139$\pm$0.009 0.102$\pm$0.011 0.099$\pm$0.021 0.094$\pm$0.015 0.126$\pm$0.013 0.193$\pm$0.012 0.123$\pm$0.014
$\Delta$ (bold - underline) +0.045 +0.056 +0.065 +0.099 +0.099 +0.080 +0.034 +0.105 +0.022 +0.129 +0.112 +0.074
Spearman Rank Correlation Coefficient
MS-Prob 0.001$\pm$0.015 0.035$\pm$0.017 0.054$\pm$0.016 0.076$\pm$0.016 0.056$\pm$0.018 0.044$\pm$0.016 0.032$\pm$0.023 0.066$\pm$0.024 0.031$\pm$0.021 0.067$\pm$0.021 0.087$\pm$0.020 0.057$\pm$0.022
PPL 0.056$\pm$0.018 0.076$\pm$0.016 0.024$\pm$0.016 0.035$\pm$0.017 -0.002$\pm$0.016 0.038$\pm$0.017 0.023$\pm$0.021 0.094$\pm$0.022 0.065$\pm$0.016 0.088$\pm$0.019 0.019$\pm$0.019 0.058$\pm$0.019
I-Emb 0.141$\pm$0.016 0.135$\pm$0.016 0.087$\pm$0.014 0.097$\pm$0.017 0.029$\pm$0.016 0.098$\pm$0.016 0.043$\pm$0.018 0.084$\pm$0.020 0.098$\pm$0.016 0.024$\pm$0.016 0.090$\pm$0.017 0.068$\pm$0.016
O-Emb 0.012$\pm$0.018 -0.098$\pm$0.016 0.076$\pm$0.016 0.054$\pm$0.017 0.035$\pm$0.017 0.016$\pm$0.017 0.010$\pm$0.019 0.059$\pm$0.025 -0.077$\pm$0.019 0.084$\pm$0.019 0.067$\pm$0.023 0.029$\pm$0.021
TV score (Ours) 0.165$\pm$0.013 0.110$\pm$0.012 0.141$\pm$0.012 0.086$\pm$0.015 0.115$\pm$0.014 0.123$\pm$0.013 0.112$\pm$0.012 0.164$\pm$0.012 0.125$\pm$0.018 0.168$\pm$0.018 0.161$\pm$0.014 0.146$\pm$0.015
w/ DiSmo (Ours) 0.173$\pm$0.011 0.208$\pm$0.010 0.072$\pm$0.015 0.145$\pm$0.019 0.109$\pm$0.016 0.141$\pm$0.014 0.134$\pm$0.015 0.127$\pm$0.016 0.189$\pm$0.017 0.167$\pm$0.017 0.155$\pm$0.016 0.154$\pm$0.016
$\Delta$ (bold - underline) +0.032 +0.073 +0.054 +0.048 +0.059 +0.043 +0.091 +0.070 +0.091 +0.080 +0.071 +0.086
Table 11: OOD Quality Estimation (GPT2-XL): Kendall’s $\tau$ and Spearman correlation between various OOD scores and benchmark quality metric binary matching. Each column shows the correlation when ID and OOD samples are merged. Underline denotes the SOTA among all baselines, and bold denotes the SOTA among our methods.
Appendix F Beyond Mathematical Reasoning
Our method has a wider range of application scenarios beyond mathematical reasoning. To verify generalizability, we choose the multiple-choice quizzing task, which has the same “pattern collapse” property as mathematical reasoning, since the output space is limited to the “ABCD” four options.
We select the MMLU dataset [12] and choose eight domains among it: high school mathematics, high school physics, high school chemistry, high school biology, high school geography, high school government and politics, high school psychology, high school statistics. We test eight rounds, each using one of the domains as the ID dataset and the remaining seven domains as the OOD datasets.
We use Llama2-7B as the training backbone, each model is trained for 3K steps and 8 batch sizes in 4-card NVIDIA Tesla V100 (2 per card). The AUROC score matrices are shown in Figure 5(a)-(e), presenting the results for TV score, input embedding, output embedding, and perplexity, respectively. In the figures, the Roman numeral I - VIII are represented as:
  • I = high school mathematics
  • II = high school physics
  • III = high school chemistry
  • IV = high school biology
  • V = high school geography
  • VI = high school government and politics
  • VII = high school psychology
  • VIII = high school statistics
Figure 5: AUROC score matrix in MMLU dataset of different OOD scores. Rows represent ID data, and columns represent OOD data.
We find that MS-Prob and PPL nearly fail on the multiple-choice task and the output embedding is not as excellent as expected, which is caused by the pattern collapse phenomenon.
Our method is comparable to the input embedding method and has very good absolute performances. For some far-shift OOD scenarios, e.g., mathematics-psychology (I-VII), physics-politics (II-VI), etc., performances of the input embedding method and our method are basically the same, and there exists a reasonable range of competing phenomena, e.g., our method performs more advantageously under physics-biology (II-IV), and the input embedding method is better under physics-geography (II-V). For some near-shift OOD scenarios, e.g., mathematics-statistics (I-VIII), where both domains essentially belong to the category of math, our method will be more well-performed, indicating that the embedding-based method produces performance degradation in fine-grained scenarios, while our method possesses stronger robustness.
Overall, our method is scalable and has greater advantages in fine-grained detection scenarios.
Appendix G Discussion: Extended Analysis
G.1 Hyperparameter Analysis: Smoothing Order $k$
Figure 6: Smoothing order $k$ analysis: $k$ ranges from $0-5$ (TV-MaDis for $k=0$). The upper part is for the OOD detection scenario and the lower part is for the OOD quality estimation scenario; the left part is for the far-shift OOD datasets and the right part is for the near-shift OOD datasets.
In the main experiments, we found that differential smoothing is not as effective as the basic TV score, with excellent results occurring on only a few datasets. Figure 6 visualizes the results for the OOD detection scenario with no smoothing ($k=0$) and smoothing order $k$ 1-5. We find that there is a significant effect of differential smoothing in two cases: (1) very good performance without smoothing, e.g., the precalculus dataset, where the AUROC results are almost close to 100%; and (2) significantly poor performance without smoothing, e.g., the AddSub dataset, where the FPR95 results on all other near-shift OOD datasets are 20 below, the results on this dataset are more than 80, when differential smoothing helps to eliminate some of the noise features and helps better detection.
In addition, for the case of $k>0$, the peak performance basically occurs in the case of $k=1$ or $2$, which indicates that when $k$ is too large, the phenomenon of over-smoothing tends to occur. Too much useful feature information is lost, leading to a decrease in detection accuracy.
G.2 Dilemmas of Embedding Representations in Reasoning.
OOD Domain
(a) Input Embedding MaDis (b) Output Embedding (w/ CoT) MaDis
algebra geometry cnt.&prob num.theory precalculus algebra geometry cnt.&prob num.theory precalculus
ID Domain algebra - 63.77 80.01 58.79 50.80 - 71.37 84.24 65.06 56.71
geometry 85.68 - 88.14 86.55 69.03 72.12 - 92.42 86.08 70.89
cnt.&prob 49.95 44.94 - 38.02 51.76 46.19 53.61 - 40.15 51.63
num.theory 66.35 78.87 50.00 - 71.93 63.02 78.76 69.85 - 67.96
precalculus 85.14 79.31 86.18 89.34 - 79.84 85.65 86.77 90.35 -
Table 12: AUROC score matrix produced after alternating the MATH dataset’s five domains as ID and OOD data measured by (a) Input Embedding Mahalanobis Distance and (b) Output Embedding (w/ CoT) Mahalanobis Distance. Darker colors represent better performances.
Prior work [46] has demonstrated that traditional training-free algorithms are ineffective in text generation scenarios apart from embedding-based methods. Nevertheless, mathematical reasoning renders embedding-based methods ineffective, as verified in our experiments. We now explore embedding-based methods in reasoning scenarios to understand their failure causes better.
We select SimCSE [11] as a powerful sentence embedding technique to represent input samples, and let the five domains in the MATH dataset alternately serve as ID and OOD data for detection. Table 12(a) presents the AUROC score matrix. We find that the accuracy of different ID-OOD pairs varies largely. For example, the accuracy is generally high when geometric or calculus domains are used as the ID data, but the accuracy is generally randomized when Cnt.&Prob is used as the ID data. This phenomenon can illustrate that as a semantic representation, embedding is currently unable to measure difficulty and digit accurately in mathematical reasoning scenarios.
G.3 Can Chain-of-Thought Address Pattern Collapse?
A straightforward approach to addressing pattern collapse in the output space is to leverage chain-of-thought (CoT) techniques [54] to expand the output space size. Likewise, we adopt the solution steps associated with each sample in the MATH dataset as the output and employ SimCSE to derive embedding representations. The experimental setup aligns with Sec.4.1, and the results are shown in Table 12(b). We note a similar phenomenon as in Table 12(a), i.e., the detection accuracy under different ID-OOD pairs varies greatly, and thus the detection randomness is more pronounced. This suggests that despite that CoT expands the output space size, the output answer is still essentially related to the difficulty and digit of mathematical reasoning, and the semantic embedding representation cannot reflect these features accurately.